### v2

# EUROPEAN ORGANIZATION FOR NUCLEAR RESEARCH (CERN)

EUROPEAN ORGANIZATION FOR NUCLEAR RESEARCH (CERN)
[Uncaptioned image] CERN-EP-2025-106 LHCb-PAPER-2024-042 18 November 2025
Coherent photoproduction of $\rho^{0},\omega$ and excited vector mesons in ultraperipheral PbPb collisions
LHCb collaborationAuthors are listed at the end of this paper.
The invariant-mass distribution for the coherent photoproduction of dipions in ultraperipheral PbPb collisions is measured using data, corresponding to an integrated luminosity of $224.6\pm 9.6\,\upmu\text{b}^{-1}$, collected by the LHCb experiment in 2018 at a nucleon-nucleon centre-of-mass energy $\sqrt{s_{\scriptscriptstyle\text{NN}}}=5.02\text{\,Te\kern-1.00006ptV}$. In the mass range from 400 to 1200 MeV, the results are consistent with previous experiments, with the spectrum dominated by the ${\rho}^{0}$} meson, which interferes with a nonresonant component, together with a smaller $\omega$ meson contribution. In an extended mass range up to 2300 MeV, models previously used do not fit the data and a consistent description requires the introduction of two resonances at masses of $1350\pm 20\text{\,Me\kern-1.00006ptV}$ and $1790\pm 20$ MeV with widths of about 300 MeV. The cross-section for each meson is measured differentially in twelve bins of rapidity from 2.05 to 4.90. The $\rho^{0}$ cross-section increases with rapidity from about 400 to 600 mb and is measured with a typical precision of 8%, while the cross-section times branching fraction for the $\omega,\rho^{\prime}$ and $\rho^{\prime\prime}$, with the statistical precision of the data, do not have a pronounced rapidity dependence and are between 0.5 and 1.5mb, with uncertainties up to 30%. A large nuclear suppression is observed for the $\rho^{0}$ meson compared to expectations based on photoproduction on the proton that use the impulse approximation. Significant suppression is also observed compared to that predicted by elastic scattering described in the Glauber approach, or with the addition of inelastic scattering in a Gribov–Glauber model.
Published in JHEP 11 (2025) 103.
© 2025 CERN for the benefit of the LHCb collaboration. CC BY 4.0 licence.
1 Introduction
Photoproduction of vector mesons on nuclei probes nuclear structure and is sensitive to nuclear shadowing and saturation effects [Brodsky:1994kf, Klein:2019qfb, Guzey:2020ntc, Accardi:2012qut]. The process can be studied in PbPb collisions at the LHC, when a photon is emitted from one lead ion, fluctuates into a virtual quark pair, and collides with the other target ion. The interaction with the target is mediated by colourless QCD propagators, Reggeons or Pomerons [Donnachie:2002en], and is shown diagrammatically in Fig. 1. Since the net colour flow in coherent photoproduction is zero, only a single vector meson is produced with a clear gap in rapidity relative to the outgoing ions. At small impact parameters, additional strong interactions between the ions are likely, thereby destroying the rapidity gap. However, by moving beyond the range of the strong force, only electromagnetic sources need to be considered. Experimentally, this is achieved in ultraperipheral collisions (UPC) [Bertulani:2005ru], which produce the distinctive signature of a single reconstructed meson and no additional particle production, with the ions remaining intact.
Figure 1: Graphical representation of $\rho$-meson photoproduction in PbPb collisions. The shaded area depicts the interaction of the $\rho$ meson with the target.
This paper presents a study of the dipion final state produced in PbPb UPC, by the LHCb collaboration. Although dominated by the ${\rho}^{0}$} meson, contributions are also present from the $\omega$ meson, nonresonant production, as well as possible excited vector mesons. Measurements of the cross-sections for coherent ${\rho}^{0}$} and $\omega$ meson production are performed. The measured ${\rho}^{0}$} cross-section, $\sigma_{{\rm PbPb}\rightarrow{\rm Pb}\,{{\rho}^{0}}\,{\rm Pb}}$, together with a calculation of the photon flux, then allows the photoproduction cross-section on the nucleus, $\sigma_{\gamma{\rm Pb}\rightarrow{{\rho}^{0}}{\rm Pb}}$, to be derived. A comparison of this result with photoproduction on the proton, $\sigma_{\gamma p\rightarrow{{\rho}^{0}}p}$, allows the nuclear suppression factor to be determined [Guzey:2020ntc, Mantysaari:2023xcu]. Naïvely, in the impulse approximation [Chew:1952fca], the ion-to-proton photoproduction ratio should scale as $A^{4/3}$ where $A$ is the number of nucleons. However, since the ${\rho}^{0}$} meson is produced before it hits the target [Ioffe:1969kf], it may undergo multiple interactions with the nucleons, which suppress the observable cross-section. This can be taken into account by the Gribov–Glauber mechanism [Gribov:1968jf, Glauber:1970jm] giving predictions [Guzey:2016piu] where the suppression depends on the total $\rho$-nucleon cross-section. The suppression for ${\rho}^{0}$} mesons is much larger than for $J\mskip-3.0mu/\mskip-2.0mu\psi$}} mesons, since the total ${\rho}^{0}$}–nucleon cross-section is about 20 mb, three order of magnitude greater than the $J\mskip-3.0mu/\mskip-2.0mu\psi$}-nucleon cross-section, which leads to a suppression of $\sigma_{\gamma{\rm Pb}\rightarrow{{\rho}^{0}}{\rm Pb}}$ compared to the impulse approximation by a factor five for the ${\rho}^{0}$, in contrast to about 10% for the $J\mskip-3.0mu/\mskip-2.0mu\psi$} meson. Suppression in excess of this would require additional mechanisms for nuclear shadowing [Guzey:2016qwo] such as saturation, which is expected to be enhanced in nuclear collisions [Accardi:2012qut] and at low values of Bjorken-$x$, the fractional momentum of the partons in the nucleon. In this respect, the forward region covered by the LHCb experiment is of particular interest since in ${\rho}^{0}$}-meson photoproduction, $x$ values down to $10^{-6}$ are probed.
The $\rho$ meson is a broad resonance decaying almost exclusively to two pions. Since its discovery in the early 1960s, there have been difficulties in measuring its mass and width. In photoproduction, the resonance is also skewed, which is attributed to sizeable interference with the nonresonant production of pion pairs. A 1974 review by Spital and Yennie [Spital:1974cx] summarises the situation, which is essentially unchanged today, in observing the resonance shape from reconstructing ${{\rho}^{0}}\rightarrow{{\pi}^{+}}{{\pi}^{-}}$ decays. The most precise measurements of the mass and width come from $e^{+}e^{-}\rightarrow{{\pi}^{+}}{{\pi}^{-}}$ data and $\tau$ decays, where discrepancies between these measurements seem to have been resolved [Jegerlehner:2011ti]. The quoted values in the PDG [PDG2024] have very small uncertainties on the mass, $m_{{\rho}^{0}}=775.26\pm 0.23\text{\,Me\kern-1.00006ptV}$, [^1] Natural units are used throughout this paper. and width, $\Gamma_{{\rho}^{0}}=147.4\pm 0.8\text{\,Me\kern-1.00006ptV}$, but there are several assumptions and model dependencies involved in this determination. The situation is far from settled with a recent paper finding a value for the mass about 20 MeV lower [Bartos:2017ils]. Consistency in the $\rho$ parameters and the dipion spectrum measured in different production mechanisms is important, not just for an understanding of light-meson spectroscopy, but also because these feed into the determination of the hadronic corrections to the anomalous muon magnetic moment, g2g-2 [Davier:2017zfy].
The $\rho$ parameters can be determined very precisely in PbPb collisions with the LHCb detector, making use of a large sample of about 20 million dipion events in UPC that has good invariant-mass resolution and is essentially free of background. This is due to the excellent tracking reconstruction and particle identification of the LHCb detector and the fact that in UPC, the transverse momentum, $p_{\mathrm{T}}$, being the Fourier transform of the impact parameter, is very small.
Although dominated by the ${\rho}^{0}$} meson and interference with nonresonantly-produced dipions, a precise measurement of the dipion spectrum allows other resonances to be identified. Interference between the ${\rho}^{0}$} and $\omega$ mesons, which was observed in AuAu collisions by the STAR collaboration [STAR:2017enh], in $ep$ collisions by the H1 collaboration [H1:2020lzc], and in ppPb collisions by the CMS collaboration [CMS:2019awk], is seen allowing the differential cross-section for $\omega$ production in UPC to be measured. The dipion spectrum also indicates the presence of other particles, likely to be excited ${\rho}^{0}$} mesons. In Ref. [ALICE:2020ugp], the ALICE collaboration reports an excess of events in the high-mass region, corresponding to a resonance mass of $1725\pm 17\text{\,Me\kern-1.00006ptV}$ and a width of $143\pm 21\text{\,Me\kern-1.00006ptV}$, as determined with a simple Gaussian model. With a similar model, the STAR collaboration [Klein:2016dtn] obtained a mass of $1653\pm 10\text{\,Me\kern-1.00006ptV}$ and a width of $164\pm 15\text{\,Me\kern-1.00006ptV}$, which they noted was compatible with the spin-3 $\rho_{3}(1690)$.
Measurements of UPC ${\rho}^{0}$}-meson production have been performed by the STAR collaboration [STAR:2007elq, STAR:2011wtm, STAR:2017enh] at nucleon-nucleon centre-of-mass energies, $\sqrt{s_{\rm NN}}$, of 62.4 and 200 GeV on gold nuclei, by the CMS collaboration in ppPb collisions at 5.02 TeV, while the ALICE collaboration performed measurements on xenon at 5.44 TeV [ALICE:2021jnv], and lead at 2.76 and 5.02 TeV [ALICE:2015nbw, ALICE:2020ugp, ALICE:2023jgu]. The STAR and ALICE data on coherent $\rho$ production in heavy-ion collisions were compared to the impulse approximation , Glauber, and Gribov–Glauber models in Ref.[Guzey:2016piu], where it was shown that the impulse approximation was about a factor four above the data, while the Glauber model was about a factor 1.5 too high. With the implementation of a Gribov–Glauber model, agreement with the data was achieved, although this used a parametrisation of H1 data that has since been improved [H1:2020lzc].
Recent discussion on nuclear suppression in coherent production has revolved around measurement of the $J\mskip-3.0mu/\mskip-2.0mu\psi$} meson, which have been performed by the STAR collaboration [STAR:2023vvb] and at the LHC by the ALICE [ALICE:2021tyx, ALICE:2021gpt, ALICE:2018oyo], CMS [CMS:2023snh] and LHCb [LHCb-PAPER-2022-012] collaborations, where the nuclear suppression factor is consistent with, or exceeds, what is expected in saturation models [Mantysaari:2023xcu]. Complementary measurements for the $\rho$ meson are welcome, since here the effect is much larger, although there are greater theoretical uncertainties when working in the nonperturbative regime.
The structure of this paper is as follows. In Sec. 2, the detector, data-taking conditions and simulation are described. The selection of the sample of dipions is detailed in Sec. 3. The lineshape measurement is presented in Sec. 4. The cross-section results are given in Sec. 5 and details of how these are used to extract the nuclear suppression factors are given in Sec. 6. Section 7 presents conclusions.
2 The LHCb detector, data collection and simulation
[…]
The data used in this analysis were collected in PbPb collisions in 2018 at a nucleon-nucleon centre-of-mass energy of 5.02 TeV, which corresponds to a total integrated luminosity of $224.6\pm 9.6\,\upmu\text{b}^{-1}$ [LHCb:2014vhh]. The data were selected by a low-multiplicity trigger at the hardware stage that required between one and twenty SPD deposits, which has a signal efficiency above 98% over most of the dipion mass and rapidity range, since only two SPD deposits are expected for the signal. At the software trigger stage, one or more tracks were required to be reconstructed in detectors located both upstream and downstream of the magnet.
The use of simulation is limited to correcting for geometrical acceptance, measuring inefficiencies that amount to about $\sim 3\%$, and constraining the yields of small backgrounds whose shapes are determined in data. The production and decay of ${\rho}^{0}$} mesons is generated by the STARlight program [Klein:2016yzr], which is also used to investigate the two-photon production of muon and electron pairs. In addition, since the kinematics of the decay are fully determined once the mass, rapidity and transverse momentum of the parent meson have been specified along with the decay angle in the rest-frame of the meson, the ROOT [Antcheva:2011zz] routine TGenPhaseSpace is used to study other particle decays. The interaction of the generated particles with the detector, and its response, are modelled using the Geant4 toolkit [GEANT4:2002zbu, Allison:2006ve] as described in Ref. [Clemencic:2011zza].
3 Selection of coherent UPC-produced dipions
Coherent UPC, without the presence of additional hadronic interactions, can be selected by ensuring the impact parameter is greater than twice the radius of the ions and thus the transverse momentum of the central system is small, $p_{\mathrm{T}}\sim 40\text{\,Me\kern-1.00006ptV}$, and the ions remaining intact, so no additional particles are produced apart from the central system. For this analysis, the experimental signature is two pions and nothing else in the detector, with the pions being approximately back-to-back in the transverse plane.
Simulation is used to determine a fiducial region in which the detector reconstruction is efficient. Both tracks must have $p_{\mathrm{T}}>100\text{\,Me\kern-1.00006ptV}$ and $2<\eta<5$, and the parent meson must have a rapidity, $y$, in the range $2.05<y<4.90$ and a mass above 400 MeV. The angle, $\theta^{*}$, between the outgoing positively charged pion and the outgoing lead ion, in the rest frame of the parent meson, must satisfy $|\cos\theta^{*}|$ below a value that varies from 0.7, between rapidities of 3 and 4, to 0.2 at the edges of the detector.
Dipion candidates are formed by requiring precisely two oppositely charged tracks reconstructed in the fiducial region with no additional tracks or photons. The tracks must form a vertex that is spatially consistent with the average interaction region. Both tracks must be consistent with the pion hypothesis created using information from the RICH detectors. The contamination from dikaons in the sample is estimated to be below 0.5% and is determined using $\phi\rightarrow K^{+}K^{-}$ decays that can be reconstructed in events where either track is consistent with the kaon hypothesis. Neither track should be identified as a muon or electron. For rapidities above 3.25, this reduces the contamination from the $\gamma\gamma\rightarrow ee$ and $\gamma\gamma\rightarrow\mu\mu$ processes to below 0.5%, while below rapidities of 3.25, the maximum contamination is 8% and concentrated at the lowest dipion masses. Low activity is required in the HeRSCheL detector, which has a good efficiency to detect nuclear breakup. To suppress multihadron contamination including $\omega\rightarrow{{\pi}^{+}}{{\pi}^{-}}{{\pi}^{0}}$ and ${{\rho}^{0}}\rightarrow{{\pi}^{+}}{{\pi}^{-}}\gamma$ decays, it is required that the transverse momentum of the dipion system is below 100 MeV, which reduces this background to below 0.5%. The final sample contains about 20 million candidates.
Acceptance and efficiency correction factors, $\varepsilon$, are determined as a function of the dipion mass, rapidity and $\cos\theta^{*}$, assuming that the $\rho$ meson inherits the transverse polarisation of the photon and decays following a $1-\cos^{2}\theta^{*}$ distribution. The acceptance is mainly a geometric effect and is determined in simulation as the fraction of mesons decaying inside the fiducial volume, taking account of the magnetic field that can bend low-momentum particles out of the active region of the detector. At central rapidities it is close to 70% decreasing to 5% in the far-forward region. A systematic uncertainty of between 4 and 6%, which is the largest single source of uncertainty, is assigned to this estimation by comparing the response of known pions from the decay ${{K}^{0}_{\mathrm{S}}}\rightarrow{{\pi}^{+}}{{\pi}^{-}}$ in data and simulation.
The ${{K}^{0}_{\mathrm{S}}}\rightarrow{{\pi}^{+}}{{\pi}^{-}}$ sample in data is also used to determine the pion-identification efficiency, which is typically $(90\pm 3)\%$ where the uncertainty arises from kinematic differences between the $\rho$ and ${K}^{0}_{\mathrm{S}}$} mesons. The efficiency due to the requirement that there are no photons reconstructed in the event is calculated per rapidity bin, using events with identified photons but a well-balanced $p_{\mathrm{T}}$ between the two pions. This is typically 94% and a systematic uncertainty of 2% is assigned as the largest difference with simulation. The vetoes on identified electrons or muons have typical efficiencies of $(98\pm 1)\%$. The efficiency for the HeRSCheL veto on activity varies from 80% to 85% with rapidity. That veto is estimated to remove about half of the events having an additional electromagnetic interaction between the nuclei [Guzey:2013jaa], which leads to excitation of one or both ions, a subsequent neutron emission (seen using zero-degree calorimeters by the ALICE and CMS [CMS:2023snh, ALICE:2023jgu] collaborations), and partial or full breakup of the nucleus. Such events are identified by the presence of a diffractive peak in the $p_{\mathrm{T}}$ distribution of the vetoed events and a systematic uncertainty of 2% is assigned, reflecting the precision with which the shape of the incoherent component is determined. Efficiencies for the other selection requirements are above 94%. The total systematic uncertainty on the efficiency of the selection is about 7% of which 4% is uncorrelated between rapidity bins.
4 Fit to the dipion spectrum
The fractions of dipions associated to the decay of the $\rho$ meson and to other processes are found from a $\chi^{2}$ fit to the number of events, $N$, binned in dipion invariant mass, $m_{\pi\pi}$. The data are fitted with a function of the form
$\int\varepsilon(m^{\prime}){\cal S}(m^{\prime},\alpha_{i})R(m_{\pi\pi},m^{\prime})\mathrm{d}m^{\prime}+{\cal B}(m_{\pi\pi}),$ (1)
[…]
For the signal function, several different parametrisations are used. First, a model, due to Söding [Soding:1965nh] that has contributions for a $\rho$ resonance and a nonresonant continuum that is constant with mass. Second, an extension to this that includes a contribution for the $\omega$ resonance. The third model uses a more flexible shape to describe the continuum. Finally, two additional resonances are introduced to describe the spectrum in an extended mass range.
The first parametrisation of the dipion spectrum is due to Söding [Soding:1965nh], which models the $\rho$ meson using a Breit–Wigner function with mass, $m_{\rho}$, and an energy-dependent width, $\Gamma$, and includes interference with a nonresonant contribution described using a constant:
${\cal S}(m_{\pi\pi})\propto\left|f_{\rho}\frac{\sqrt{m_{\pi\pi}m_{\rho}\Gamma(m_{\pi\pi})}}{m_{\pi\pi}^{2}-m_{\rho}^{2}+im_{\rho}\Gamma(m_{\pi\pi})}+f_{\rm nr}\right|^{2},$~ (2)
with
$\Gamma(m_{\pi\pi})=\Gamma_{\rho}\frac{m_{\rho}}{m_{\pi\pi}}\frac{q(m_{\pi\pi})}{q(m_{\rho})},$ (3)
where $\Gamma_{\rho}$ is the pole width, $q(m)=\sqrt{m^{2}-4m_{\pi}^{2}}$ is the momentum of the pion with mass $m_{\pi}$ in the centre-of-mass of a system with mass $m$, while $f_{\rho}$ and $f_{\rm nr}$ are free parameters, describing the $\rho$ and nonresonant contributions, respectively.
Table 1: Results of the fit to LHCb data that uses Eq. 2, compared to measurements by other experiments. The uncertainties on the LHCb values are dominated by systematic effects.
Experiment LHCb ZEUS [ZEUS:1997rof] H1 [H1:2009cml, H1:1996prv] ALICE [ALICE:2015nbw]
Fit range [ GeV ] [0.4,1.2] [0.55,1.20] [0.6,1.1] [0.6,1.5]
$m_{\rho}$ [ MeV ] $71\pm 3 77$$70\pm 2$ $769\pm 4$ $762.0^{+6.5}_{-3.8}$
$\Gamma_{\rho}$ [ MeV ] $149\pm 4$ $146\pm 13$ $162\pm 8$ $150^{+13}_{-7}$
$|f_{\rm nr}/f_{\rho}|$ [ $\text{\,Ge\kern-1.00006ptV}^{-{1}/{2}}]$ $0.72\pm 0.04$ $0.70\pm 0.04$ $0.57\pm 0.09$ $0.50^{+0.10}_{-0.06}$
The fit is performed for $m_{\pi\pi}$ between 400 MeV and 1200 MeV in twelve bins of rapidity. Compared to measurements at other experiments [ZEUS:1997rof, H1:1996prv, H1:2009cml, H1:2020lzc, STAR:2007elq, STAR:2011wtm, STAR:2017enh, CMS:2019awk, ALICE:2015nbw, ALICE:2020ugp], it is possible to extend the fit to lower mass values due to a reduced background contamination and the ability to reconstruct tracks down to a $p_{\mathrm{T}}$ of 100 MeV. The results of the fit for the rapidity range $2.75<y<4.50$ are given in Table 1 and are similar to those from other experiments for the mass and width, while the nonresonant contribution differs between experiments, most likely due to different amounts of residual background. Due to the large signal sample, the statistical uncertainties on the fit parameters are below 0.1%. A systematic uncertainty is estimated from the standard deviation of the results in the individual rapidity bins and by fixing the background contribution to zero.
Figure 2: Dipion invariant-mass distribution with the result of the fit superimposed, where the signal is modelled by the function in Eq. 2. The lower panel shows the normalised residuals.
Figure 3: Invariant-mass fits to data where the signal model includes a contribution for the $\omega$ resonance with the parametrisation of (left) Eq. 4 and (right) Eq. 5.
The fit is shown in Fig. 2. Although the overall shape is similar to the data, the fit quality is poor, particularly in the region of the $\omega$ resonance. Similar observations were made by the STAR [STAR:2011wtm], H1 [H1:2020lzc], and CMS [CMS:2019awk] collaborations and led to the introduction of a second Breit–Wigner function. Following the STAR parametrisation, the signal function is extended to read
${\cal S}(m_{\pi\pi})=\left|f_{\rho}\frac{\sqrt{m_{\pi\pi}m_{\rho}\Gamma(m_{\pi\pi})}}{m_{\pi\pi}^{2}-m_{\rho}^{2}+im_{\rho}\Gamma(m_{\pi\pi})}+f_{\rm nr}+f_{\omega}\exp(i\phi_{\omega})\frac{\sqrt{m_{\pi\pi}m_{\omega}\Gamma_{\omega}}}{m_{\pi\pi}^{2}-m_{\omega}^{2}+im_{\omega}\Gamma_{\omega}}\right|^{2},$~ (4)
where $f_{\omega}$ is the contribution of the $\omega$ resonance and $\phi_{\omega}$ is its phase relative to the $\rho$ meson. The mass, $m_{\omega}$, and width, $\Gamma_{\omega}$, are fixed to the PDG values [PDG2024] and a constant width is assumed for $\Gamma_{\omega}$ as it is very narrow: implementing an energy-dependent width for the $\omega$ resonance makes minimal difference to the fit results. The left panel of Fig. 3 shows that the quality of the fit improves considerably in the peak region, and although the absolute size of the $\omega$ contribution is small, its presence and the resulting $\rho-\omega$ interference has a significant impact. The parameters of the fit are presented in Table 2 and are consistent with the STAR results.
Table 2: Results of fits to LHCb data that include an $\omega$ resonance using the parametrisation of Eq. 4, compared to measurements by other experiments. The uncertainties quoted for the LHCb measurements are the estimated systematic contributions.
LHCb STAR [STAR:2017enh] CMS [CMS:2019awk]
$m_{\rho}$ [ MeV ] 774 $\pm$ 3 776.2 $\pm$ 0.2 773 $\pm$ 1
$\Gamma_{\rho}$ [ MeV ] 153 $\pm$ 3 148 $\pm$ 1 148 $\pm$ 3
$|f_{\rm nr}/f_{\rho}|[\text{\,Ge\kern-1.00006ptV}^{-1/2}]$ 0.73 $\pm$ 0.03 0.79 $\pm$ 0.08 0.50 $\pm$ 0.06
$\phi_{\omega}$ [rad] 1.4 $\pm$ 0.1 1.46 $\pm$ 0.11 1.8 $\pm$ 0.3
$|f_{\omega}/f_{\rho}|$ 0.32 $\pm$ 0.03 0.36 $\pm$ 0.05 0.40 $\pm$ 0.06
Table 3: Results of fits to LHCb data that include an $\omega$ resonance using the parametrisation of Eq. 5, compared to the measurement by the H1 collaboration. The uncertainties quoted for the LHCb measurements are the estimated systematic contributions.
LHCb H1 [H1:2020lzc]
$m_{\rho}$ [ MeV ] 774 $\pm$ 3 770.8 $\pm$ $1.3^{+2.3}_{-2.4}$
$\Gamma_{\rho}$ [ MeV ] 153 $\pm$ 3 151.3 $\pm$ $2.2^{+1.6}_{-2.8}$
$|f_{\rm nr}/f_{\rho}|$ 0.18 $\pm$ 0.02 0.189 $\pm$ $0.026^{+0.025}_{-0.016}$
[…]
$|f_{\omega}/f_{\rho}|$ 0.15 $\pm$ 0.01 0.166 $\pm$ $0.017^{+0.008}_{-0.023}$
$\Lambda$ [ MeV ] 400 $\pm$ 100 180 $\pm$ $590^{+200}_{-100}$
$\delta$ 1.1 $\pm$ 0.1 0.76 $\pm$ $0.35^{+0.14}_{-0.08}$
The H1 collaboration used a different parametrisation of the phase and interference and also included two additional parameters to allow the nonresonant shape to differ from a constant:
${\cal S}(m_{\pi\pi})=\frac{q^{3}(m_{\pi\pi})}{q^{3}(m_{\rho})}\left|{f_{\rho}\,\cal{BW}}_{\rho}(m_{\pi\pi})\biggl(1+f_{\omega}\exp{({i\phi_{\omega}})}\frac{m_{\pi\pi}^{2}}{m_{\omega}^{2}}{\cal{BW}}_{\omega}(m_{\pi\pi})\biggr)+\frac{f_{\rm nr}}{(m_{\pi\pi}^{2}-4m_{\pi}^{2}+\Lambda^{2})^{\delta}}\right|^{2}\,$,~ (5)
where
${\cal{BW}}_{V}(m_{\pi\pi})=\frac{m_{V}\Gamma_{V}}{m_{\pi\pi}^{2}-m_{V}^{2}+im_{V}\Gamma(m_{\pi\pi})}\,$, (6)
are Breit–Wigner distributions for the $\rho$ and $\omega$ resonances (with $V=\rho,\omega)$, and $\Lambda$ and $\delta$ are free parameters. Adopting their convention gives a better fit to the data as can be seen in the right panel of Fig. 3. The results are summarised in Table 3 and are consistent with the H1 results.
Figure 4: Invariant-mass fits to data (left) in the region between 400 and 1200 MeV, where the signal contribution is parametrised by Eq. 5, and (right) in the full region, where the signal model is parametrised by Eq. 7, which includes two additional resonances and a phase for the nonresonant component.
Although the fit is acceptable in the fitted region, when expanding the fit range up to dipion masses of 2.3 GeV, displayed in the left panel of Fig. 4, the model overshoots the data by an order of magnitude. Furthermore, the data show a clear structure at around 1.7 GeV, which is also discernible in STAR [Klein:2016dtn] and ALICE [ALICE:2020ugp] data. To improve the quality of the fit to the extended mass range, two modifications are made to the signal model that now takes the form:
[…]
First, a phase factor, $\phi_{\rm nr}$, is introduced between the $\rho$ and nonresonant components and this leads to a reduction of the signal function at higher masses. Second, to describe the bump around 1.7 GeV and the cusp at about 1.2 GeV, two extra Breit–Wigner resonances are introduced to describe excited $\rho$ states, designated as $\rho^{\prime}$ and $\rho^{\prime\prime}$, with $f_{\rho\prime}$ and $f_{\rho\prime\prime}$ the contributions of each. The parameters of the fit are given in Table 4 and the result is shown in the right panel of Fig. 4. Excellent agreement with data is obtained across the full range.
There is a long history of possible vector states in the high-mass region as discussed in the PDG review [PDG2024]. The $\rho^{\prime}$ resonance is measured to have a mass of $1350\pm 5\pm 20\text{\,Me\kern-1.00006ptV}$ and a width of $320\pm 10\pm 40\text{\,Me\kern-1.00006ptV}$, while the $\rho^{\prime\prime}$ resonance has a mass of $1790\pm 5\pm 20\text{\,Me\kern-1.00006ptV}$ and a width of $290\pm 10\pm 40\text{\,Me\kern-1.00006ptV}$, where the first uncertainty is the statistical precision of the fit result and the second is due to systematic effects, evaluated using the standard deviation of results obtained by fitting the model separately in each rapidity bin. The statistical significance of either or both resonances compared to the null hypothesis of Eq. 5 greatly exceeds the threshold for claiming an observation. However, it should be pointed out that improvements to the fit have been obtained with alternative parametrisations including using a form factor to multiply $\cal{BW}_{\rho}$, and including a contribution from a noninterfering $f_{2}(1270)$ meson.
These measurements are consistent with the values of the $\rho(1450)$ and $\rho(1700)$ mesons in the PDG [PDG2024], noting the role the choice of dynamic lineshapes can play. Further evidence for the existence of these particles was reported by the Babar collaboration using $e^{+}e^{-}\rightarrow\pi\pi(\gamma)$ data [BaBar:2012bdw] where the same qualitative features are observed as in the LHCb data. The production processes are similar with the main difference being that in photoproduction the photon is quasi-real while it is strongly virtual in $e^{+}e^{-}$ annihilation. The coupling between the photon and the pions is probed in both processes so the LHCb data may help clarify the resonant contributions, which have an impact in determining the hadronic corrections to the g2g-2 measurement [Davier:2019can].
Table 4: Result of the fit to data with the signal model of Eq. 7, which includes a phase for the nonresonant component, and Breit–Wigner functions for the $\omega$, $\rho^{\prime}$ and $\rho^{\prime\prime}$ resonances. The uncertainties are dominated by systematic effects.
Parameter (Eq. 7) Result
$m_{\rho}$ [ MeV ] 774 $\pm$ 3
$\Gamma_{\rho}$ [ MeV ] 148 $\pm$ 7
$|f_{\rm nr}/f_{\rho}|$ 0.18 $\pm$ 0.03
$\phi_{\rm nr}$ 0.1-$\pm 0.1$
$\Lambda$ [ MeV ] 420 $\pm$ 150
$\delta$ 1.2 $\pm$ 0.1
$|f_{\omega}/f_{\rho}|$ 0.14 $\pm$ 0.02
$\phi_{\omega}$ [ rad ] $20±\pm 0.0$6
$|f_{\rho\prime}/f_{\rho}|$ 0.013 $\pm$ 0.008
$m_{\rho\prime}$ [ MeV ] 1350 $\pm$ 20
$\Gamma_{\rho\prime}$ [ MeV ] 320 $\pm$ 40
$\phi_{\rho^{\prime}}$ 2.4 $\pm$ 0.3
$|f_{\rho\prime\prime}/f_{\rho}|$ 0.006 $\pm$ 0.002
$m_{\rho\prime\prime}$ [ MeV ] 1790 $\pm$ 20
$\Gamma_{\rho\prime\prime}$ [ MeV ] 290 $\pm$ 40
$\phi_{\rho^{\prime\prime}}$ 1.2 $\pm$ 0.7
5 Cross-section
The production cross-section, $\sigma_{V}$, for a vector meson, $V$, is determined in twelve bins of rapidity from 2.05 to 4.90 through
$\frac{\mathrm{d}\sigma_{V}}{\mathrm{d}y}=\frac{f_{\rm smog}N_{V}}{\Delta_{y}\,{\cal L}\,{\rm BF}_{V\rightarrow\pi\pi}},$ (8)
[…]
The number of mesons in each rapidity bin is given by
$N_{V}=\int_{2m_{\pi}}^{m_{V}+5\Gamma_{V}}{\cal S}(m)\mathrm{d}m\,$, (9)
with $\cal S$, depending on the model, defined in Eq. 45 or 7, with the coefficients $f_{\rho},f_{\omega},f_{\rho\prime},f_{\rho\prime\prime},f_{\rm nr}$ set to zero except for the meson of interest. In order to compare the cross-section results across experiments, the integration is performed, by convention, from threshold to $m_{V}+5\Gamma_{V}$, which are taken to be 1510, 826, 2950 and 3240 MeV for the $\rho,\omega,\rho^{\prime}$ and $\rho^{\prime\prime}$ meson, respectively. The reference fits for the measurement of the $\rho$ and $\omega$ cross-section use the model of Eq. 5. Fitting with the model of Eq. 4 increases the $\rho$ cross-section on average by 2% but increases that of the $\omega$ meson by 34%, showing the model-dependence of the results. The fit with the model of Eq. 7 increases the $\rho$ cross-section by 6%, while that of the $\omega$ meson is unchanged. The total systematic uncertainty for the $\rho$ cross-section varies from 7 to 10%, of which about 4% is uncorrelated between bins.
To improve the stability of the $\omega$ cross-section as a function of rapidity, the fit range is constrained to be between 600 and 900  MeV, and a correlated systematic uncertainty is assigned corresponding to the standard deviation of differences to the fit result performed between 400 and 1200 MeV. This is the dominant uncertainty. The fits to extract the cross-section for the $\rho^{\prime}$ and $\rho^{\prime\prime}$ are performed in the mass range from 1.0 to 2.3 GeV, with all parameters fixed except for $f_{\rho\prime}$ and $f_{\rho\prime\prime}$. The uncertainty includes a correlated systematic component corresponding to the standard deviation of differences compared to the fit with all parameters free in the full mass range. Also in this case, it is the dominant uncertainty.
Table 5: Differential cross-section for the $\rho,\omega,\rho^{\prime}$ and $\rho^{\prime\prime}$ mesons in bins of rapidity, using the model of Eq. 7. The first uncertainty includes statistical and uncorrelated systematics. The second uncertainty represents the correlated systematics.
Rapidity $[2.05,2.25]$ $[2.25,2.50]$ $[2.50,2.75]$ $[2.75,3.00]$
$\frac{\mathrm{d}\sigma_{\rho}}{\mathrm{d}y}$[mb] 473 $\pm$ 23  $\pm$ 40 412 $\pm$ 16  $\pm$ 33 434 $\pm$ 17  $\pm$ 34 453 $\pm$ 18  $\pm$ 36
$\frac{\mathrm{d}\sigma_{\omega}}{\mathrm{d}y}$[mb] 52 $\pm$ 10  $\pm$ 6 24 $\pm$ 4  $\pm$ 3 31 $\pm$ 3  $\pm$ 4 32 $\pm$ 4  $\pm$ 4
${\rm BF}(\rho^{\prime}\rightarrow\pi\pi)\frac{\mathrm{d}\sigma_{\rho}\prime}{\mathrm{d}y}$[mb] 1.55 $\pm$ 0.12  $\pm$ 0.33 1.36 $\pm$ 0.08  $\pm$ 0.29 1.46 $\pm$ 0.07  $\pm$ 0.31 1.46 $\pm$ 0.07  $\pm$ 0.31
${\rm BF}(\rho^{\prime\prime}\rightarrow\pi\pi)\frac{\mathrm{d}\sigma_{\rho}\prime\prime}{\mathrm{d}y}$[mb] 0.89 $\pm$ 0.07  $\pm$ 0.28 1.04 $\pm$ 0.05  $\pm$ 0.32 1.15 $\pm$ 0.06  $\pm$ 0.36 1.26 $\pm$ 0.06  $\pm$ 0.39
Rapidity $[3.00,3.25]$ $[3.25,3.50]$ $[3.50,3.75]$ $[3.75,4.00]$
$\frac{\mathrm{d}\sigma_{\rho}}{\mathrm{d}y}$[mb] 464 $\pm$ 18  $\pm$ 37 477 $\pm$ 19  $\pm$ 33 487 $\pm$ 19  $\pm$ 33 493 $\pm$ 19  $\pm$ 33
$\frac{\mathrm{d}\sigma_{\omega}}{\mathrm{d}y}$[mb] 42 $\pm$ 3  $\pm$ 5 45 $\pm$ 3  $\pm$ 5 45 $\pm$ 3  $\pm$ 5 36 $\pm$ 3  $\pm$ 4
${\rm BF}(\rho^{\prime}\rightarrow\pi\pi)\frac{\mathrm{d}\sigma_{\rho}\prime}{\mathrm{d}y}$[mb] 1.33 $\pm$ 0.07  $\pm$ 0.29 1.18 $\pm$ 0.59  $\pm$ 0.25 1.03 $\pm$ 0.05  $\pm$ 0.22 0.93 $\pm$ 0.05  $\pm$ 0.20
${\rm BF}(\rho^{\prime\prime}\rightarrow\pi\pi)\frac{\mathrm{d}\sigma_{\rho}\prime\prime}{\mathrm{d}y}$[mb] 1.35 $\pm$ 0.06  $\pm$ 0.42 1.22 $\pm$ 0.05  $\pm$ 0.38 1.08 $\pm$ 0.05  $\pm$ 0.33 0.92 $\pm$ 0.05  $\pm$ 0.28
Rapidity $[4.00,4.25]$ $[4.25,4.50]$ $[4.50,4.75]$ $[4.75,4.90]$
$\frac{\mathrm{d}\sigma_{\rho}}{\mathrm{d}y}$[mb] 498 $\pm$ 20  $\pm$ 28 520 $\pm$ 21  $\pm$ 30 556 $\pm$ 22  $\pm$ 38 625 $\pm$ 43  $\pm$ 44
$\frac{\mathrm{d}\sigma_{\omega}}{\mathrm{d}y}$[mb] 43 $\pm$ 3  $\pm$ 4 46 $\pm$ 4  $\pm$ 5 55 $\pm$ 6  $\pm$ 6 41 $\pm$ 14  $\pm$ 4
${\rm BF}(\rho^{\prime}\rightarrow\pi\pi)\frac{\mathrm{d}\sigma_{\rho}\prime}{\mathrm{d}y}$[mb] 1.08 $\pm$ 0.06  $\pm$ 0.22 1.29 $\pm$ 0.06  $\pm$ 0.27 1.44 $\pm$ 0.10  $\pm$ 0.30 1.56 $\pm$ 0.29  $\pm$ 0.33
${\rm BF}(\rho^{\prime\prime}\rightarrow\pi\pi)\frac{\mathrm{d}\sigma_{\rho}\prime\prime}{\mathrm{d}y}$[mb] 0.60 $\pm$ 0.03  $\pm$ 0.18 0.56 $\pm$ 0.03  $\pm$ 0.17 0.29 $\pm$ 0.03  $\pm$ 0.09 0.02 $\pm$ 0.04  $\pm$ 0.01
Figure 5: Differential cross-section for coherent $\rho$ photoproduction as a function of rapidity. The LHCb points use the model of Eq. 7 and the error bars represent the statistical and total uncertainties. The shaded band corresponds to the estimated uncertainty on the GKZ prediction. The contributions due to the photon emitted from each ion are labelled as $W^{+}$ and $W^{-}$ and are explained further in Sec. 6. Also shown are ALICE results at central rapidity and the output of the STARlight generator.
Figure 6: Differential cross-section multiplied by the branching fraction to dipions, in bins of rapidity, for $\omega,\rho^{\prime}$ and $\rho^{\prime\prime}$ coherent photoproduction, calculated using the model of Eq. 7. The error bars represent the statistical uncertainty, while the bands correspond to the total uncertainty.
The results are given in Table 5 with the uncorrelated and correlated uncertainties quoted separately, and are shown in Figs. 5 and  6. For the $\rho$ meson, a comparison is made to ALICE data [ALICE:2020ugp], to the predictions of Guzey, Kryshen and Zhalov (GKZ) [Guzey:2016piu], and to the results of the STARlight generator [Klein:2016yzr]. Further details of this calculation are provided in Sec. 6. The trend in the data for $\rho$ production is reproduced by the GKZ model, which takes account of elastic and inelastic nuclear shadowing. However, the GKZ model predicts a slightly higher overall cross-section. In other words, the data prefer more suppression than is given by the QCD effects considered in Ref. [Guzey:2016piu]. The measured cross-section roughly agrees with the STARlight generator although the $\rho$ measurement is systematically higher for rapidities above 3.5. However, no strong conclusions can be drawn since STARlight simply uses a nuclear form factor to scale the HERA photoproduction results (see the comment in Sec. 4 of Ref.[Frankfurt:2015cwa]).
The results for the $\omega$ photoproduction cross-section have a strong fit-model dependence with the ratio $\sigma_{\rho}/\sigma_{\omega}$ varying between $9.8\pm 1.0$ (Eq. 4), $12.9\pm 1.3$ (Eq. 5) and $13.5\pm 1.4$ (Eq. 7). These results are in agreement with the determinations of this ratio by the ZEUS [ZEUS:1996zse] and H1 [H1:2020lzc] collaborations of $8.6\pm 1.0$ and $10.3^{+3.3}_{-2.3}$, respectively, using photoproduction on the proton at similar photon-proton centre-of-mass energies. The ZEUS collaboration measured the $\omega$ cross-section directly in the channel $\omega\rightarrow{{\pi}^{+}}{{\pi}^{-}}{{\pi}^{0}}$, while H1 measured it through its interference, as performed here, in the channel $\omega\rightarrow{{\pi}^{+}}{{\pi}^{-}}$.
6 Nuclear suppression factor
Nuclear suppression effects are accessed by comparing the scaling between photoproduction on the proton, $\sigma_{\gamma p\rightarrow{{\rho}^{0}}p}$, and photoproduction on the nucleus, $\sigma_{\gamma{\rm Pb}\rightarrow{{\rho}^{0}}{\rm Pb}}$. First, the $\sigma_{\gamma p\rightarrow{{\rho}^{0}}p}$ results from the H1 [H1:2020lzc] experiment are scaled to $\sigma_{\gamma{\rm Pb}\rightarrow{{\rho}^{0}}{\rm Pb}}$ using three models described below: the impulse approximation, a Glauber model, and a Gribov–Glauber model. Then, after accounting for the photon fluxes from the lead nuclei, a prediction is obtained for the coherent production of $\rho$ mesons in PbPb collisions, which is compared to the experimental data.
The differential cross-section for ${\rm PbPb}\rightarrow{\rm Pb}\rho{\rm Pb}$ can be expressed [Guzey:2016piu] in terms of the photoproduction cross-section $\gamma{\rm Pb}\rightarrow\rho{\rm Pb}$ using the photon flux $\mathrm{d}N_{\gamma}/\mathrm{d}k$ for the number of photons $N_{\gamma}$ of energy $k$:
$\frac{\mathrm{d}\sigma_{{\rm PbPb}\rightarrow{\rm Pb}\,\rho\,{\rm Pb}}}{\mathrm{d}y}=\biggl(k\frac{\mathrm{d}N_{\gamma}}{\mathrm{d}k}\biggr)^{+}\sigma_{\gamma{\rm Pb}\rightarrow\rho{\rm Pb}}(W^{+})+\biggl(k\frac{\mathrm{d}N_{\gamma}}{\mathrm{d}k}\biggr)^{-}\sigma_{\gamma{\rm Pb}\rightarrow\rho{\rm Pb}}(W^{-}).$ (10)
The two terms on the right-hand side correspond to each of the ions being the photon emitter. For LHCb, the ‘+’ index corresponds to the higher-energy photon that is emitted by the ion moving with positive rapidity towards the detector. The photon energy, $k^{\pm}=(m_{\rho}/2)\exp(\pm y)$, and the centre-of-mass energy of the photon-ion system, $W^{\pm}=(2k^{\pm}\sqrt{s})^{0.5}$, are related to the fractional momentum of the partons in the nucleon by $x=m_{\rho}^{2}/W^{2}$. Thus for PbPb UPC in LHCb, $W^{+}$ is in the range 170–710 GeV corresponding to $x\sim[1,19]\times 10^{-6}$ where Pomeron exchange dominates, while $W^{-}$ is in the range 5–22 GeV, and $x\sim[0.001,0.020]$, where Reggeon exchange is also important.
In the absence of nuclear effects, the $\rho$ photoproduction cross-section on the ion scales from the cross-section on the proton according to the number of nucleons and is given by the impulse approximation [Chew:1952fca]. However, due to the large cross-section for $\rho$-nucleon scattering, as the meson traverses the nucleus, multiple interactions are possible. This can be taken into account in the Glauber model that considers elastic scatters, or extended to include inelastic scatters with a Gribov–Glauber model [Frankfurt:2015cwa]. Details of these calculations are given in the Appendix. The predictions for the impulse approximation, Glauber and Gribov–Glauber models are shown in the left panel of Fig. 7, where it can be seen that the elastic and inelastic rescattering effects have a major impact on the result. The data are clearly inconsistent with the impulse approximation, while nuclear shadowing alone is also insufficient to provide a complete description. Including inelastic scattering improves the overall picture but some additional suppression is needed.
Figure 7: Differential cross-section for the $\rho$ meson as a function of rapidity for data (left) compared to the impulse approximation, Glauber model, and Gribov–Glauber model and (right) with the estimated contributions from photoproduction on each ion in the Gribov–Glauber model after including nuclear suppression factors. The impulse approximation has been scaled down by a factor of two so it can be viewed with this axis-range.
Equation 10 shows there is a two-fold ambiguity in determining the photon emitter. However, at the highest rapidities, irrespective of nuclear suppression, the backward-going lead nucleus is almost always the photon emitter. As a result, the first term on the right-hand side only contributes 3% in the rapidity bin from 4.75 to 4.90, well within the systematic uncertainties. Ignoring this ‘+’ solution, Eq. 10 predicts a $\rho$ cross-section for PbPb collisions without nuclear suppression of $3330\pm 770$ mb, to be compared with the measured value of $625\pm 43\pm 44$ mb. In the perturbative regime where the Pomeron is equivalent to two gluons, the nuclear suppression factor has been defined [Guzey:2020ntc] with respect to the impulse approximation using
$S_{\rm IA}=\sqrt{\frac{\sigma_{\gamma\rm Pb\rightarrow\rho Pb}}{\sigma^{\rm IA}_{\gamma\rm Pb\rightarrow\rho Pb}}},$ (11)
since for perturbatively produced objects like the $J\mskip-3.0mu/\mskip-2.0mu\psi$}} meson, this quantity is related to the ratio of the gluon parton distribution functions in the nucleus and the proton. The same definition of $S$ is used here, although the relationship with the parton distribution functions is not possible using the nonperturbatively produced $\rho$ meson. A suppression factor of $S_{\rm IA}=0.43\pm 0.05$ is measured at a centre-of-mass energy $W=5.4\text{\,Ge\kern-1.00006ptV}$ or equivalently, a fractional parton momentum of $x=0.018$.
The LHCb data also have sensitivity to the nuclear suppression factor at high $W$ (low $x$), since at low rapidities the ‘+’ solution contributes about one third. A fit to all the data is performed as follows. Equation 10 is modified to include a nuclear suppression factor,
$\frac{\mathrm{d}\sigma_{{\rm PbPb}\rightarrow{\rm Pb}\,\rho\,{\rm Pb}}}{\mathrm{d}y}=(S^{+}_{\rm model})^{2}\biggl(k\frac{\mathrm{d}N_{\gamma}}{\mathrm{d}k}\biggr)^{+}\sigma^{\rm model}_{\gamma{\rm Pb}\rightarrow\rho{\rm Pb}}(W^{+})+(S^{-}_{\rm model})^{2}\biggl(k\frac{\mathrm{d}N_{\gamma}}{\mathrm{d}k}\biggr)^{-}\sigma^{\rm model}_{\gamma{\rm Pb}\rightarrow\rho{\rm Pb}}(W^{-}),$ (12)
where one suppression factor is used for high and one for low $W$, and
$S_{\rm model}=\sqrt{\frac{\sigma_{\gamma\rm Pb\rightarrow\rho Pb}}{\sigma^{\rm model}_{\gamma\rm Pb\rightarrow\rho Pb}}},$ (13)
is more-generally defined with respect to the impulse approximation, Glauber, or Gribov-Glauber model.
A chisquare fit is performed with a Lagrange multiplier to take account of the correlated uncertainties, assuming that the uncertainties on the theory predictions are all correlated. The quantity that is minimised is
$\chi^{2}(S^{+},S^{-},\lambda)=\sum_{i}\biggl(\frac{x^{\rm exp}_{i}-x^{\rm th}_{i}}{\sigma^{\rm uncor}_{i}}\biggr)^{2}+\lambda^{2}\,$, (14)
where $x^{\rm exp}$ are the measured values in Table 5, $\sigma_{i}^{\rm uncor}$ are the uncorrelated experimental uncertainties, $x^{\rm th}_{i}={\mathrm{d}\sigma_{\rm PbPb\rightarrow Pb\rho Pb}/\mathrm{d}y}+\lambda\sigma_{i}^{\rm cor}$ with $\mathrm{d}\sigma_{\rm PbPb\rightarrow Pb\rho Pb}/\mathrm{d}y$} given by Eq. 12, and $\sigma_{i}^{\rm cor}$ the correlated uncertainties associated with the experimental uncertainties, the photon flux, the parametrisation of $\sigma_{\rho N}$, and the dispersion of $P(\sigma)$ (as described in the Appendix).
With the suppression factor measured relative to the impulse approximation, the fit returns values of $S_{\rm IA}^{-}=0.39\pm 0.03$ and $S_{\rm IA}^{+}=0.05\pm 0.11$ while $\lambda=0.6\pm 0.6$. This result shows that the events produced by higher-energy photons (lower-energy gluons) experience more nuclear suppression than those produced by lower-energy photons (higher energy gluons), as predicted by the Glauber and Gribov–Glauber calculations. A similar effect, albeit at a much reduced level in the pertubative $J\mskip-3.0mu/\mskip-2.0mu\psi$} process, has been reported in Refs. [Guzey:2020ntc, Mantysaari:2023xcu] using ALICE and CMS UPC data [ALICE:2023jgu, CMS:2023snh]. To take account of rescattering contributions, the nuclear suppression is calculated with respect to the Glauber approximation that includes elastic contributions predicted by QCD. This returns $S_{\rm Glauber}^{-}=0.78\pm 0.04$ and $S_{\rm Glauber}^{+}=0.30\pm 0.11$ while $\lambda=0.9\pm 0.9$. This shows that nonlinear effects are much more significant for the low-$x$ gluons. Finally, after including inelastic contributions using the Gribov–Glauber model, the relative suppression factors are $S_{\rm Gribov-Glauber}^{-}=0.84\pm 0.04$ and $S^{+}_{\rm Gribov-Glauber}=0.83\pm 0.06$ while $\lambda=0.7\pm 0.9$. The result is shown in the right panel of Fig. 7. The data is in closer agreement to the model although some additional suppression is still needed, which could be due to saturation effects or an increased dispersion of $P(\sigma)$.
7 Conclusions
Using a clean sample of dipions produced in ultraperipheral PbPb collisions, the invariant-mass distribution shows a prominent ${\rho}^{0}$} peak, but also contributions from nonresonant production, the $\omega$ resonance and higher-mass resonances. Hints of a resonance at 1.7 GeV reported by the ALICE and STAR collaborations are unambiguously confirmed with the larger LHCb data sample that, in the model chosen, also requires a resonance at 1.35 GeV. The cross-sections for coherent photoproduction of ${{\rho}^{0}},\omega,\rho^{\prime}$ and $\rho^{\prime\prime}$ mesons are measured as functions of rapidity. The coherent $\omega$ cross-section is found to be about a factor ten lower than that of the ${\rho}^{0}$} meson in agreement with photoproduction measurements at HERA.
The differential cross-section for the ${\rho}^{0}$} meson shows a slow rise with rapidity, in agreement with the GKZ model that includes linear and nonlinear QCD effects using a Gribov–Glauber calculation to describe multiple nucleon interactions. The observed cross-section is lower by about 30%, indicating that additional suppression is required. In PbPb interactions, there is a two-fold ambiguity in determining the photon emitter. However, by utilising the different cross-section dependencies with rapidity, there is sensitivity to both high and low $x$-values. Compared to the Gribov–Glauber model, suppression factors of $0.84\pm 0.04$ for values of $x\sim 5\times 10^{-3}$ and of $0.83\pm 0.06$ for $x\sim 5\times 10^{-6}$ are obtained. Additional suppression is observed and this could be caused by saturation effects. However, models that explicitly include saturation cannot be used to predict the cross-section since the ${\rho}^{0}$} meson is too light to apply perturbation theory. We thus rely on a phenomenological model that scales and extrapolates fits to fixed-target and H1 data for photoproduction on the proton. The LHCb photon-nucleon centre-of-mass energies in the range 170–760 GeV require an extrapolation of the H1 data, while energies in the range 5–23 GeV are only measured in fixed-target experiments, are less precise, and are also in a region where Reggeon exchange is important. Thus, new measurements of photoproduction on the proton are highly desirable at these energies and can be obtained at forward rapidities in proton-lead collisions at the LHC, or using fixed-target gaseous collisions at LHCb.
Appendix
Using the photon flux, Eq. 10 relates the differential cross-section for ${\rm PbPb}\rightarrow{\rm Pb}\rho{\rm Pb}$ to the photoproduction cross-section $\gamma{\rm Pb}\rightarrow\rho{\rm Pb}$, expressed using the impulse approximation, the Glauber model, or the Gribov–Glauber model.
The photon flux for quasi-real photons can be calculated in the equivalent photon approximation [Baur:2001jj] as a function of $k$ and impact parameter $b$. Restricting $b$ to be greater than twice the radius of the ions, as is required in UPC, gives
$k\frac{\mathrm{d}N_{\gamma}}{\mathrm{d}k}=\frac{2Z^{2}\alpha}{\pi}\biggl(\zeta K_{0}(\zeta)K_{1}(\zeta)-\frac{\zeta^{2}}{2}[K_{1}^{2}(\zeta)-K^{2}_{0}(\zeta)]\biggr),$ (15)
with $\zeta=2R_{\rm Pb}k/\upgamma$, where $R_{\rm Pb}=6.68\text{\,fm}$ [Jones:2014aoa] is the radius of the ion, $\upgamma$ is the nucleus Lorentz factor, $Z$ is the charge of the ion, $\alpha$ is the fine-structure constant and $K_{0},K_{1}$ are modified Bessel functions.
The impulse approximation is given by
$\sigma^{\rm IA}_{\gamma{\rm Pb}\rightarrow\rho{\rm Pb}}(W)=\frac{\mathrm{d}\sigma_{\gamma p\rightarrow\rho p}}{\mathrm{d}t}(W,t=0)\int_{-\infty}^{t_{\rm min}}|F_{\rm Pb}(t)|^{2}\mathrm{d}t,$ (16)
which only takes into account the nuclear form factor, $F_{\rm Pb}(t)$, and neglects all other potential nuclear effects. The four-momentum-transfer squared is represented by $t\approx-p_{\mathrm{T}}^{2}$ with $t_{\rm min}=-m_{\rho}^{4}m_{p}^{2}/{W^{4}}$. The form factor is a Fourier transform of the nuclear density,
[…]
The H1 collaboration measured $\rho$ production in the energy range $20<W<80\text{\,Ge\kern-1.00006ptV}$, while the range considered here is larger. For $W<20\text{\,Ge\kern-1.00006ptV}$, Reggeon as well as Pomeron contributions are important, while for $W>80\text{\,Ge\kern-1.00006ptV}$ an extrapolation of the HERA data is required. By fitting their data combined with fixed target data, the H1 collaboration found [H1:2020lzc]
$\sigma_{\gamma p\rightarrow\rho p}(W)=\sigma(W_{0})\biggl\{\biggl(\frac{W}{W_{0}}\biggr)^{\delta_{\cal P}}+f_{\cal R}\biggl(\frac{W}{W_{0}}\biggr)^{\delta_{\cal R}}\biggr\}\,$, (19)
with a cross-section of $\sigma(W_{0})=10.66\pm 0.11^{+0.75}_{-1.00}\,\upmu\text{b}$ at $W_{0}=40\text{\,Ge\kern-1.00006ptV}$, while the Pomeron and Reggeon powers are $\delta_{\cal P}=0.207\pm 0.015^{+0.053}_{-0.033}$ and $\delta_{\cal R}=-1.45\pm 0.12^{+0.35}_{-0.21}$, and the fraction of Reggeon contribution, $f_{\cal R}=0.020\pm 0.007^{+0.029}_{-0.013}$. The $t$-dependence for their data is approximately exponential: $\frac{\mathrm{d}\sigma}{\mathrm{d}t}=\frac{\mathrm{d}\sigma}{\mathrm{d}t}|_{t=0}\exp(bt)$, which by integration yields $\frac{\mathrm{d}\sigma}{\mathrm{d}t}|_{t=0}=b\sigma_{\gamma p\rightarrow\rho p}$. The $b$-dependence was also determined with H1 and fixed-target data using just a single equation of the form
$b(W)=b(W_{0})+4\alpha_{1}\log(W/W_{0})\,$, (20)
that describes the data well in both the Reggeon- and Pomeron-dominated regions. The shrinkage of the diffraction peak was observed with fitted values of $b(W_{0})=9.59\pm 0.14^{+0.14}_{-0.12}\text{\,Ge\kern-1.00006ptV}^{-2}$ and $\alpha_{1}=0.233\pm 0.064^{+0.020}_{-0.038}\text{\,Ge\kern-1.00006ptV}^{-2}$. Combining these results gives
$\frac{\mathrm{d}\sigma_{\gamma p\rightarrow\rho p}}{\mathrm{d}t}(W,t=0)=\biggl(b(W_{0})+4\alpha_{1}\log(W/W_{0})\biggr)\sigma(W_{0})\biggl\{\biggl(\frac{W}{W_{0}}\biggr)^{\delta_{\cal P}}+f_{\cal R}\biggl(\frac{W}{W_{0}}\biggr)^{\delta_{\cal R}}\biggr\}.$ (21)
Using this expression in Eq. 16 and Eq. 10 allows a calculation of the differential cross-section for UPC $\rho$ production in the impulse approximation.
Multiple interactions are taken into account using a Glauber calculation, assuming vector-meson dominance and using the optical theorem to relate the total cross-section to the imaginary part of the forward scattering amplitude. Following Ref. [Frankfurt:2015cwa], the photoproduction cross-section on the nucleus can be written,
$\sigma^{\rm Glauber}_{\gamma{\rm Pb}\rightarrow\rho{\rm Pb}}=\biggl(\frac{e}{f_{\rho}}\biggr)^{2}\int\mathrm{d}^{2}\vec{b}\,\biggl|1-e^{-\sigma_{\rho N}T_{A}(b)/2}\biggr|^{2}\,$, (22)
where $T_{A}$ is the thickness of the nucleus at impact parameter $\vec{b}$ and $\sigma_{\rho N}$ is the total $\rho$–nucleon cross-section. The experimentally derived quantity $f_{\rho}^{2}/(4\pi)=2.01\pm 0.01$ is used for the $\rho$$\gamma$ coupling constant. The thickness is found from the nuclear density,
[…]
The total cross-section for $\rho$–nucleon scattering can be found once again from the optical theorem,
$\frac{\mathrm{d}\sigma_{\gamma p\rightarrow\rho p}}{\mathrm{d}t}(W,t=0)=\frac{1}{16\pi}\biggl(\frac{e}{f_{\rho}}\biggr)^{2}\sigma^{2}_{\rho N}(W),$ (24)
with $\mathrm{d}\sigma_{\gamma p\rightarrow\rho p}/{\mathrm{d}t}$ given by Eq. 21.
The Glauber approximation only considers elastic scatters and needs to be extended to include inelastic scatters. This can be done with a Gribov–Glauber model as described in Ref. [Frankfurt:2015cwa], and Eq. 22 is modified to
$\sigma^{\rm Gribov-Glauber}_{\gamma{\rm Pb}\rightarrow\rho{\rm Pb}}=\biggl(\frac{e}{f_{\rho}}\biggr)^{2}\int\mathrm{d}^{2}\vec{b}\,\biggl|\int\mathrm{d}\sigma P(\sigma)\biggl(1-e^{-\sigma T_{A}(b)/2}\bigg)\biggr|^{2},$ (25)
where now instead of a single value for $\sigma_{\rho N}$, a distribution, $P(\sigma)$, of $\rho$-nucleon cross-sections is used with
$P(\sigma)=C\frac{1}{1+(\sigma/\sigma_{0})^{2}}e^{-(\sigma/\sigma_{0}-1)^{2}/\Omega^{2}}.$ (26)
The parameters, $C,\sigma_{0},\Omega$ are detailed in Ref. [Frankfurt:2015cwa], where $C$ is a normalisation constant, the average total cross-section $\int\sigma P(\sigma)\mathrm{d}\sigma=\sigma_{\rho N}$, and the width of the distribution, $\Omega$ is constrained by data.
The uncertainties on the predictions are calculated as the quadratic sum of the uncertainty on the photon flux, obtained by changing the ion radius by 0.5 fm, and the uncertainty on $\mathrm{d}\sigma_{\gamma p\rightarrow p}/\mathrm{d}t(W,t=0)$, calculated from the uncertainties and correlations on the parameters of Eq. 21 given in Ref. [H1:2020lzc]. For the Gribov–Glauber model, an additional uncertainty is added in quadrature, obtained by changing the dispersion of $P(\sigma)$ as described in Ref. [Frankfurt:2015cwa].
Acknowledgements
We express our gratitude to our colleagues in the CERN accelerator departments for the excellent performance of the LHC. We thank the technical and administrative staff at the LHCb institutes. We acknowledge support from CERN and from the national agencies: ARC (Australia); CAPES, CNPq, FAPERJ and FINEP (Brazil); MOST and NSFC (China); CNRS/IN2P3 (France); BMBF, DFG and MPG (Germany); INFN (Italy); NWO (Netherlands); MNiSW and NCN (Poland); MCID/IFA (Romania); MICIU and AEI (Spain); SNSF and SER (Switzerland); NASU (Ukraine); STFC (United Kingdom); DOE NP and NSF (USA). We acknowledge the computing resources that are provided by ARDC (Australia), CBPF (Brazil), CERN, IHEP and LZU (China), IN2P3 (France), KIT and DESY (Germany), INFN (Italy), SURF (Netherlands), Polish WLCG (Poland), IFIN-HH (Romania), PIC (Spain), CSCS (Switzerland), and GridPP (United Kingdom). We are indebted to the communities behind the multiple open-source software packages on which we depend. Individual groups or members have received support from Key Research Program of Frontier Sciences of CAS, CAS PIFI, CAS CCEPP, Fundamental Research Funds for the Central Universities, and Sci. & Tech. Program of Guangzhou (China); Minciencias (Colombia); EPLANET, Marie Skłodowska-Curie Actions, ERC and NextGenerationEU (European Union); A*MIDEX, ANR, IPhU and Labex P2IO, and Région Auvergne-Rhône-Alpes (France); Alexander-von-Humboldt Foundation (Germany); ICSC (Italy); Severo Ochoa and María de Maeztu Units of Excellence, GVA, XuntaGal, GENCAT, InTalent-Inditex and Prog. Atracción Talento CM (Spain); SRC (Sweden); the Leverhulme Trust, the Royal Society and UKRI (United Kingdom).
We thank Misha Ryskin for many helpful discussions and Vadim Guzey for his assistance with the calculation of the nuclear suppression factors and for making his results available to us.
LHCb collaboration
R. Aaij38[Uncaptioned image], A.S.W. Abdelmotteleb57[Uncaptioned image], C. Abellan Beteta51[Uncaptioned image], F. Abudinén57[Uncaptioned image], T. Ackernley61[Uncaptioned image], A. A.  Adefisoye69[Uncaptioned image], B. Adeva47[Uncaptioned image], M. Adinolfi55[Uncaptioned image], P. Adlarson82[Uncaptioned image], C. Agapopoulou14[Uncaptioned image], C.A. Aidala83[Uncaptioned image], Z. Ajaltouni11, S. Akar66[Uncaptioned image], K. Akiba38[Uncaptioned image], P. Albicocco28[Uncaptioned image], J. Albrecht19,f[Uncaptioned image], F. Alessio49[Uncaptioned image], Z. Aliouche63[Uncaptioned image], P. Alvarez Cartelle56[Uncaptioned image], R. Amalric16[Uncaptioned image], S. Amato3[Uncaptioned image], J.L. Amey55[Uncaptioned image], Y. Amhis14[Uncaptioned image], L. An6[Uncaptioned image], L. Anderlini27[Uncaptioned image], M. Andersson51[Uncaptioned image], A. Andreianov44[Uncaptioned image], P. Andreola51[Uncaptioned image], M. Andreotti26[Uncaptioned image], D. Andreou69[Uncaptioned image], A. Anelli31,o[Uncaptioned image], D. Ao7[Uncaptioned image], F. Archilli37,v[Uncaptioned image], M. Argenton26[Uncaptioned image], S. Arguedas Cuendis9,49[Uncaptioned image], A. Artamonov44[Uncaptioned image], M. Artuso69[Uncaptioned image], E. Aslanides13[Uncaptioned image], R. Ataíde Da Silva50[Uncaptioned image], M. Atzeni65[Uncaptioned image], B. Audurier12[Uncaptioned image], D. Bacher64[Uncaptioned image], I. Bachiller Perea10[Uncaptioned image], S. Bachmann22[Uncaptioned image], M. Bachmayer50[Uncaptioned image], $J.$ Back57[Uncaptioned image], P. Baladron Rodriguez47[Uncaptioned image], V. Balagura15[Uncaptioned image], A.  Balboni26[Uncaptioned image], W. Baldini26[Uncaptioned image], L. Balzani19[Uncaptioned image], H.  Bao7[Uncaptioned image], J. Baptista de Souza Leite61[Uncaptioned image], C. Barbero Pretel47,12[Uncaptioned image], M. Barbetti27[Uncaptioned image], I. R. Barbosa70[Uncaptioned image], R.J. Barlow63[Uncaptioned image], M. Barnyakov25[Uncaptioned image], S. Barsuk14[Uncaptioned image], W. Barter59[Uncaptioned image], M. Bartolini56[Uncaptioned image], J. Bartz69[Uncaptioned image], J.M. Basels17[Uncaptioned image], S. Bashir40[Uncaptioned image], G. Bassi35,s[Uncaptioned image], B. Batsukh5[Uncaptioned image], P. B.  Battista14[Uncaptioned image], A. Bay50[Uncaptioned image], A. Beck57[Uncaptioned image], M. Becker19[Uncaptioned image], F. Bedeschi35[Uncaptioned image], I.B. Bediaga2[Uncaptioned image], N. A.  Behling19[Uncaptioned image], S. Belin47[Uncaptioned image], K. Belous44[Uncaptioned image], I. Belov29[Uncaptioned image], I. Belyaev36[Uncaptioned image], G. Benane13[Uncaptioned image], G. Bencivenni28[Uncaptioned image], E. Ben-Haim16[Uncaptioned image], A. Berezhnoy44[Uncaptioned image], R. Bernet51[Uncaptioned image], S. Bernet Andres46[Uncaptioned image], A. Bertolin33[Uncaptioned image], C. Betancourt51[Uncaptioned image], F. Betti59[Uncaptioned image], J.  Bex56[Uncaptioned image], Ia. Bezshyiko51[Uncaptioned image], J. Bhom41[Uncaptioned image], M.S. Bieker19[Uncaptioned image], N.V. Biesuz26[Uncaptioned image], P. Billoir16[Uncaptioned image], A. Biolchini38[Uncaptioned image], M. Birch62[Uncaptioned image], F.C.R. Bishop10[Uncaptioned image], A. Bitadze63[Uncaptioned image], A. Bizzeti27,p[Uncaptioned image], T. Blake57[Uncaptioned image], F. Blanc50[Uncaptioned image], J.E. Blank19[Uncaptioned image], S. Blusk69[Uncaptioned image], V. Bocharnikov44[Uncaptioned image], J.A. Boelhauve19[Uncaptioned image], O. Boente Garcia15[Uncaptioned image], T. Boettcher66[Uncaptioned image], A.  Bohare59[Uncaptioned image], A. Boldyrev44[Uncaptioned image], C.S. Bolognani79[Uncaptioned image], R. Bolzonella26,l[Uncaptioned image], R. B.  Bonacci1[Uncaptioned image], N. Bondar44[Uncaptioned image], A. Bordelius49[Uncaptioned image], F. Borgato33,q[Uncaptioned image], S. Borghi63[Uncaptioned image], M. Borsato31,o[Uncaptioned image], J.T. Borsuk41[Uncaptioned image], S.A. Bouchiba50[Uncaptioned image], M.  Bovill64[Uncaptioned image], T.J.V. Bowcock61[Uncaptioned image], A. Boyer49[Uncaptioned image], C. Bozzi26[Uncaptioned image], A. Brea Rodriguez50[Uncaptioned image], N. Breer19[Uncaptioned image], J. Brodzicka41[Uncaptioned image], A. Brossa Gonzalo47,†[Uncaptioned image], J. Brown61[Uncaptioned image], D. Brundu32[Uncaptioned image], E. Buchanan59[Uncaptioned image], A. Buonaura51[Uncaptioned image], L. Buonincontri33,q[Uncaptioned image], A.T. Burke63[Uncaptioned image], C. Burr49[Uncaptioned image], J.S. Butter56[Uncaptioned image], J. Buytaert49[Uncaptioned image], W. Byczynski49[Uncaptioned image], S. Cadeddu32[Uncaptioned image], H. Cai74[Uncaptioned image], A. Caillet16[Uncaptioned image], R. Calabrese26,l[Uncaptioned image], S. Calderon Ramirez9[Uncaptioned image], L. Calefice45[Uncaptioned image], S. Cali28[Uncaptioned image], M. Calvi31,o[Uncaptioned image], M. Calvo Gomez46[Uncaptioned image], P. Camargo Magalhaes2,z[Uncaptioned image], J. I. Cambon Bouzas47[Uncaptioned image], P. Campana28[Uncaptioned image], D.H. Campora Perez79[Uncaptioned image], A.F. Campoverde Quezada7[Uncaptioned image], S. Capelli31[Uncaptioned image], L. Capriotti26[Uncaptioned image], R. Caravaca-Mora9[Uncaptioned image], A. Carbone25,j[Uncaptioned image], L. Carcedo Salgado47[Uncaptioned image], R. Cardinale29,m[Uncaptioned image], A. Cardini32[Uncaptioned image], P. Carniti31,o[Uncaptioned image], L. Carus22[Uncaptioned image], A. Casais Vidal65[Uncaptioned image], R. Caspary22[Uncaptioned image], G. Casse61[Uncaptioned image], M. Cattaneo49[Uncaptioned image], G. Cavallero26,49[Uncaptioned image], V. Cavallini26,l[Uncaptioned image], S. Celani22[Uncaptioned image], D. Cervenkov64[Uncaptioned image], S.  Cesare30,n[Uncaptioned image], A.J. Chadwick61[Uncaptioned image], I. Chahrour83[Uncaptioned image], M. Charles16[Uncaptioned image], Ph. Charpentier49[Uncaptioned image], E.  Chatzianagnostou38[Uncaptioned image], M. Chefdeville10[Uncaptioned image], C. Chen13[Uncaptioned image], S. Chen5[Uncaptioned image], Z. Chen7[Uncaptioned image], A. Chernov41[Uncaptioned image], S. Chernyshenko53[Uncaptioned image], X.  Chiotopoulos79[Uncaptioned image], V. Chobanova81[Uncaptioned image], S. Cholak50[Uncaptioned image], M. Chrzaszcz41[Uncaptioned image], A. Chubykin44[Uncaptioned image], V. Chulikov28[Uncaptioned image], P. Ciambrone28[Uncaptioned image], X. Cid Vidal47[Uncaptioned image], G. Ciezarek49[Uncaptioned image], P. Cifra49[Uncaptioned image], P.E.L. Clarke59[Uncaptioned image], M. Clemencic49[Uncaptioned image], H.V. Cliff56[Uncaptioned image], J. Closier49[Uncaptioned image], C. Cocha Toapaxi22[Uncaptioned image], V. Coco49[Uncaptioned image], J. Cogan13[Uncaptioned image], E. Cogneras11[Uncaptioned image], L. Cojocariu43[Uncaptioned image], S.  Collaviti50[Uncaptioned image], P. Collins49[Uncaptioned image], T. Colombo49[Uncaptioned image], M. Colonna19[Uncaptioned image], A. Comerma-Montells45[Uncaptioned image], L. Congedo24[Uncaptioned image], A. Contu32[Uncaptioned image], N. Cooke60[Uncaptioned image], I. Corredoira 47[Uncaptioned image], A. Correia16[Uncaptioned image], G. Corti49[Uncaptioned image], J. Cottee Meldrum55[Uncaptioned image], B. Couturier49[Uncaptioned image], D.C. Craik51[Uncaptioned image], M. Cruz Torres2,g[Uncaptioned image], E. Curras Rivera50[Uncaptioned image], R. Currie59[Uncaptioned image], C.L. Da Silva68[Uncaptioned image], S. Dadabaev44[Uncaptioned image], L. Dai71[Uncaptioned image], X. Dai6[Uncaptioned image], E. Dall’Occo49[Uncaptioned image], J. Dalseno47[Uncaptioned image], C. D’Ambrosio49[Uncaptioned image], J. Daniel11[Uncaptioned image], A. Danilina44[Uncaptioned image], P. d’Argent24[Uncaptioned image], G. Darze3[Uncaptioned image], A.  Davidson57[Uncaptioned image], J.E. Davies63[Uncaptioned image], A. Davis63[Uncaptioned image], O. De Aguiar Francisco63[Uncaptioned image], C. De Angelis32,k[Uncaptioned image], F. De Benedetti49[Uncaptioned image], J. de Boer38[Uncaptioned image], K. De Bruyn78[Uncaptioned image], S. De Capua63[Uncaptioned image], M. De Cian22[Uncaptioned image], U. De Freitas Carneiro Da Graca2,a[Uncaptioned image], E. De Lucia28[Uncaptioned image], J.M. De Miranda2[Uncaptioned image], L. De Paula3[Uncaptioned image], M. De Serio24,h[Uncaptioned image], P. De Simone28[Uncaptioned image], F. De Vellis19[Uncaptioned image], J.A. de Vries79[Uncaptioned image], F. Debernardis24[Uncaptioned image], D. Decamp10[Uncaptioned image], V. Dedu13[Uncaptioned image], S.  Dekkers1[Uncaptioned image], L. Del Buono16[Uncaptioned image], B. Delaney65[Uncaptioned image], H.-P. Dembinski19[Uncaptioned image], J. Deng8[Uncaptioned image], V. Denysenko51[Uncaptioned image], O. Deschamps11[Uncaptioned image], F. Dettori32,k[Uncaptioned image], B. Dey77[Uncaptioned image], P. Di Nezza28[Uncaptioned image], I. Diachkov44[Uncaptioned image], S. Didenko44[Uncaptioned image], S. Ding69[Uncaptioned image], L. Dittmann22[Uncaptioned image], V. Dobishuk53[Uncaptioned image], A. D.  Docheva60[Uncaptioned image], C. Dong4,b[Uncaptioned image], A.M. Donohoe23[Uncaptioned image], F. Dordei32[Uncaptioned image], A.C. dos Reis2[Uncaptioned image], A. D.  Dowling69[Uncaptioned image], W. Duan72[Uncaptioned image], P. Duda80[Uncaptioned image], M.W. Dudek41[Uncaptioned image], L. Dufour49[Uncaptioned image], V. Duk34[Uncaptioned image], P. Durante49[Uncaptioned image], M. M. Duras80[Uncaptioned image], J.M. Durham68[Uncaptioned image], O. D.  Durmus77[Uncaptioned image], A. Dziurda41[Uncaptioned image], A. Dzyuba44[Uncaptioned image], S. Easo58[Uncaptioned image], E. Eckstein18[Uncaptioned image], U. Egede1[Uncaptioned image], A. Egorychev44[Uncaptioned image], V. Egorychev44[Uncaptioned image], S. Eisenhardt59[Uncaptioned image], E. Ejopu63[Uncaptioned image], L. Eklund82[Uncaptioned image], M. Elashri66[Uncaptioned image], J. Ellbracht19[Uncaptioned image], S. Ely62[Uncaptioned image], A. Ene43[Uncaptioned image], J. Eschle69[Uncaptioned image], S. Esen22[Uncaptioned image], T. Evans63[Uncaptioned image], F. Fabiano32,k[Uncaptioned image], L.N. Falcao2[Uncaptioned image], Y. Fan7[Uncaptioned image], B. Fang7[Uncaptioned image], L. Fantini34,r,49[Uncaptioned image], M. Faria50[Uncaptioned image], K.  Farmer59[Uncaptioned image], D. Fazzini31,o[Uncaptioned image], L. Felkowski80[Uncaptioned image], M. Feng5,7[Uncaptioned image], M. Feo19[Uncaptioned image], A. Fernandez Casani48[Uncaptioned image], M. Fernandez Gomez47[Uncaptioned image], A.D. Fernez67[Uncaptioned image], F. Ferrari25,j[Uncaptioned image], F. Ferreira Rodrigues3[Uncaptioned image], M. Ferrillo51[Uncaptioned image], M. Ferro-Luzzi49[Uncaptioned image], S. Filippov44[Uncaptioned image], R.A. Fini24[Uncaptioned image], M. Fiorini26,l[Uncaptioned image], M. Firlej40[Uncaptioned image], K.L. Fischer64[Uncaptioned image], D.S. Fitzgerald83[Uncaptioned image], C. Fitzpatrick63[Uncaptioned image], T. Fiutowski40[Uncaptioned image], F. Fleuret15[Uncaptioned image], M. Fontana25[Uncaptioned image], L. F.  Foreman63[Uncaptioned image], R. Forty49[Uncaptioned image], D. Foulds-Holt56[Uncaptioned image], V. Franco Lima3[Uncaptioned image], M. Franco Sevilla67[Uncaptioned image], M. Frank49[Uncaptioned image], E. Franzoso26,l[Uncaptioned image], G. Frau63[Uncaptioned image], C. Frei49[Uncaptioned image], D.A. Friday63[Uncaptioned image], J. Fu7[Uncaptioned image], Q. Führing19,f,56[Uncaptioned image], Y. Fujii1[Uncaptioned image], T. Fulghesu16[Uncaptioned image], E. Gabriel38[Uncaptioned image], G. Galati24[Uncaptioned image], M.D. Galati38[Uncaptioned image], A. Gallas Torreira47[Uncaptioned image], D. Galli25,j[Uncaptioned image], S. Gambetta59[Uncaptioned image], M. Gandelman3[Uncaptioned image], P. Gandini30[Uncaptioned image], B.  Ganie63[Uncaptioned image], H. Gao7[Uncaptioned image], R. Gao64[Uncaptioned image], T.Q. Gao56[Uncaptioned image], Y. Gao8[Uncaptioned image], Y. Gao6[Uncaptioned image], Y. Gao8[Uncaptioned image], L.M. Garcia Martin50[Uncaptioned image], P. Garcia Moreno45[Uncaptioned image], J. García Pardiñas49[Uncaptioned image], P.  Gardner67[Uncaptioned image], K. G.  Garg8[Uncaptioned image], L. Garrido45[Uncaptioned image], C. Gaspar49[Uncaptioned image], R.E. Geertsema38[Uncaptioned image], $L.$ Gerken19[Uncaptioned image], E. Gersabeck63[Uncaptioned image], M. Gersabeck20[Uncaptioned image], T. Gershon57[Uncaptioned image], S. Ghizzo29,m[Uncaptioned image], Z. Ghorbanimoghaddam55[Uncaptioned image], L. Giambastiani33,q[Uncaptioned image], F. I. Giasemis16,e[Uncaptioned image], V. Gibson56[Uncaptioned image], H.K. Giemza42[Uncaptioned image], A.L. Gilman64[Uncaptioned image], M. Giovannetti28[Uncaptioned image], A. Gioventù45[Uncaptioned image], L. Girardey63,58[Uncaptioned image], P. Gironella Gironell45[Uncaptioned image], C. Giugliano26,l[Uncaptioned image], M.A. Giza41[Uncaptioned image], E.L. Gkougkousis62[Uncaptioned image], F.C. Glaser14,22[Uncaptioned image], $V.$ Gligorov16,49[Uncaptioned image], C. Göbel70[Uncaptioned image], E. Golobardes46[Uncaptioned image], D. Golubkov44[Uncaptioned image], A. Golutvin62,49,44[Uncaptioned image], S. Gomez Fernandez45[Uncaptioned image], W.  Gomulka40[Uncaptioned image], F. Goncalves Abrantes64[Uncaptioned image], M. Goncerz41[Uncaptioned image], G. Gong4,b[Uncaptioned image], J. A. Gooding19[Uncaptioned image], I.V. Gorelov44[Uncaptioned image], C. Gotti31[Uncaptioned image], J.P. Grabowski18[Uncaptioned image], L.A. Granado Cardoso49[Uncaptioned image], E. Graugés45[Uncaptioned image], E. Graverini50,t[Uncaptioned image], L. Grazette57[Uncaptioned image], G. Graziani27[Uncaptioned image], A. T. Grecu43[Uncaptioned image], L.M. Greeven38[Uncaptioned image], N.A. Grieser66[Uncaptioned image], L. Grillo60[Uncaptioned image], S. Gromov44[Uncaptioned image], C.  Gu15[Uncaptioned image], M. Guarise26[Uncaptioned image], L.  Guerry11[Uncaptioned image], M. Guittiere14[Uncaptioned image], V. Guliaeva44[Uncaptioned image], P. A. Günther22[Uncaptioned image], A.-K. Guseinov50[Uncaptioned image], E. Gushchin44[Uncaptioned image], Y. Guz6,49,44[Uncaptioned image], T. Gys49[Uncaptioned image], K. Habermann18[Uncaptioned image], T. Hadavizadeh1[Uncaptioned image], C. Hadjivasiliou67[Uncaptioned image], G. Haefeli50[Uncaptioned image], C. Haen49[Uncaptioned image], M. Hajheidari49, G.  Hallett57[Uncaptioned image], $M.$ Halvorsen49[Uncaptioned image], P.M. Hamilton67[Uncaptioned image], J. Hammerich61[Uncaptioned image], Q. Han8[Uncaptioned image], X. Han22,49[Uncaptioned image], S. Hansmann-Menzemer22[Uncaptioned image], L. Hao7[Uncaptioned image], N. Harnew64[Uncaptioned image], T. H.  Harris1[Uncaptioned image], M. Hartmann14[Uncaptioned image], S. Hashmi40[Uncaptioned image], J. He7,c[Uncaptioned image], F. Hemmer49[Uncaptioned image], C. Henderson66[Uncaptioned image], R.D.L. Henderson1,57[Uncaptioned image], A.M. Hennequin49[Uncaptioned image], K. Hennessy61[Uncaptioned image], L. Henry50[Uncaptioned image], J. Herd62[Uncaptioned image], P. Herrero Gascon22[Uncaptioned image], J. Heuel17[Uncaptioned image], A. Hicheur3[Uncaptioned image], G. Hijano Mendizabal51[Uncaptioned image], J. Horswill63[Uncaptioned image], R. Hou8[Uncaptioned image], Y. Hou11[Uncaptioned image], N. Howarth61[Uncaptioned image], J. Hu72[Uncaptioned image], W. Hu6[Uncaptioned image], X. Hu4,b[Uncaptioned image], W. Huang7[Uncaptioned image], W. Hulsbergen38[Uncaptioned image], R.J. Hunter57[Uncaptioned image], M. Hushchyn44[Uncaptioned image], D. Hutchcroft61[Uncaptioned image], M. Idzik40[Uncaptioned image], D. Ilin44[Uncaptioned image], P. Ilten66[Uncaptioned image], A. Inglessi44[Uncaptioned image], A. Iniukhin44[Uncaptioned image], A. Ishteev44[Uncaptioned image], K. Ivshin44[Uncaptioned image], R. Jacobsson49[Uncaptioned image], H. Jage17[Uncaptioned image], S.J. Jaimes Elles75,49,48[Uncaptioned image], S. Jakobsen49[Uncaptioned image], E. Jans38[Uncaptioned image], B.K. Jashal48[Uncaptioned image], A. Jawahery67,49[Uncaptioned image], V. Jevtic19,f[Uncaptioned image], E. Jiang67[Uncaptioned image], X. Jiang5,7[Uncaptioned image], Y. Jiang7[Uncaptioned image], Y. J.  Jiang6[Uncaptioned image], M. John64[Uncaptioned image], A.  John Rubesh Rajan23[Uncaptioned image], D. Johnson54[Uncaptioned image], C.R. Jones56[Uncaptioned image], T.P. Jones57[Uncaptioned image], S. Joshi42[Uncaptioned image], B. Jost49[Uncaptioned image], J.  Juan Castella56[Uncaptioned image], N. Jurik49[Uncaptioned image], I. Juszczak41[Uncaptioned image], D. Kaminaris50[Uncaptioned image], S. Kandybei52[Uncaptioned image], M.  Kane59[Uncaptioned image], Y. Kang4,b[Uncaptioned image], C. Kar11[Uncaptioned image], M. Karacson49[Uncaptioned image], D. Karpenkov44[Uncaptioned image], A. Kauniskangas50[Uncaptioned image], J.W. Kautz66[Uncaptioned image], M.K. Kazanecki41[Uncaptioned image], F. Keizer49[Uncaptioned image], M. Kenzie56[Uncaptioned image], T. Ketel38[Uncaptioned image], B. Khanji69[Uncaptioned image], A. Kharisova44[Uncaptioned image], S. Kholodenko35,49[Uncaptioned image], G. Khreich14[Uncaptioned image], T. Kirn17[Uncaptioned image], V.S. Kirsebom31,o[Uncaptioned image], O. Kitouni65[Uncaptioned image], S. Klaver39[Uncaptioned image], N. Kleijne35,s[Uncaptioned image], K. Klimaszewski42[Uncaptioned image], M.R. Kmiec42[Uncaptioned image], S. Koliiev53[Uncaptioned image], L. Kolk19[Uncaptioned image], A. Konoplyannikov44[Uncaptioned image], P. Kopciewicz40,49[Uncaptioned image], P. Koppenburg38[Uncaptioned image], M. Korolev44[Uncaptioned image], I. Kostiuk38[Uncaptioned image], O. Kot53[Uncaptioned image], S. Kotriakhova[Uncaptioned image], A. Kozachuk44[Uncaptioned image], P. Kravchenko44[Uncaptioned image], L. Kravchuk44[Uncaptioned image], M. Kreps57[Uncaptioned image], P. Krokovny44[Uncaptioned image], W. Krupa69[Uncaptioned image], W. Krzemien42[Uncaptioned image], O. Kshyvanskyi53[Uncaptioned image], S. Kubis80[Uncaptioned image], M. Kucharczyk41[Uncaptioned image], V. Kudryavtsev44[Uncaptioned image], E. Kulikova44[Uncaptioned image], A. Kupsc82[Uncaptioned image], B. Kutsenko13[Uncaptioned image], D. Lacarrere49[Uncaptioned image], P.  Laguarta Gonzalez45[Uncaptioned image], A. Lai32[Uncaptioned image], A. Lampis32[Uncaptioned image], D. Lancierini56[Uncaptioned image], C. Landesa Gomez47[Uncaptioned image], $J.$ Lane1[Uncaptioned image], R. Lane55[Uncaptioned image], G. Lanfranchi28[Uncaptioned image], C. Langenbruch22[Uncaptioned image], J. Langer19[Uncaptioned image], O. Lantwin44[Uncaptioned image], T. Latham57[Uncaptioned image], F. Lazzari35,t[Uncaptioned image], C. Lazzeroni54[Uncaptioned image], R. Le Gac13[Uncaptioned image], H.  Lee61[Uncaptioned image], R. Lefèvre11[Uncaptioned image], A. Leflat44[Uncaptioned image], S. Legotin44[Uncaptioned image], M. Lehuraux57[Uncaptioned image], E. Lemos Cid49[Uncaptioned image], O. Leroy13[Uncaptioned image], T. Lesiak41[Uncaptioned image], E. D. Lesser49[Uncaptioned image], B. Leverington22[Uncaptioned image], A. Li4,b[Uncaptioned image], C.  Li13[Uncaptioned image], H. Li72[Uncaptioned image], K. Li8[Uncaptioned image], L. Li63[Uncaptioned image], M. Li8[Uncaptioned image], P. Li7[Uncaptioned image], P.-R. Li73[Uncaptioned image], Q.  Li5,7[Uncaptioned image], S. Li8[Uncaptioned image], T. Li5,d[Uncaptioned image], T. Li72[Uncaptioned image], Y. Li8[Uncaptioned image], Y. Li5[Uncaptioned image], Z. Lian4,b[Uncaptioned image], X. Liang69[Uncaptioned image], S. Libralon48[Uncaptioned image], C. Lin7[Uncaptioned image], T. Lin58[Uncaptioned image], R. Lindner49[Uncaptioned image], H.  Linton62[Uncaptioned image], V. Lisovskyi50[Uncaptioned image], R. Litvinov32,49[Uncaptioned image], F. L.  Liu1[Uncaptioned image], G. Liu72[Uncaptioned image], K. Liu73[Uncaptioned image], S. Liu5,7[Uncaptioned image], W.  Liu8[Uncaptioned image], Y. Liu59[Uncaptioned image], Y. Liu73[Uncaptioned image], Y. L.  Liu62[Uncaptioned image], A. Lobo Salvia45[Uncaptioned image], A. Loi32[Uncaptioned image], T. Long56[Uncaptioned image], J.H. Lopes3[Uncaptioned image], A. Lopez Huertas45[Uncaptioned image], S. López Soliño47[Uncaptioned image], Q. Lu15[Uncaptioned image], C. Lucarelli27[Uncaptioned image], D. Lucchesi33,q[Uncaptioned image], M. Lucio Martinez79[Uncaptioned image], V. Lukashenko38,53[Uncaptioned image], Y. Luo6[Uncaptioned image], A. Lupato33,i[Uncaptioned image], E. Luppi26,l[Uncaptioned image], K. Lynch23[Uncaptioned image], X.-R. Lyu7[Uncaptioned image], G. M.  Ma4,b[Uncaptioned image], S. Maccolini19[Uncaptioned image], F. Machefert14[Uncaptioned image], F. Maciuc43[Uncaptioned image], B.  Mack69[Uncaptioned image], I. Mackay64[Uncaptioned image], L. M.  Mackey69[Uncaptioned image], L.R. Madhan Mohan56[Uncaptioned image], M. J.  Madurai54[Uncaptioned image], A. Maevskiy44[Uncaptioned image], D. Magdalinski38[Uncaptioned image], D. Maisuzenko44[Uncaptioned image], M.W. Majewski40, $J.$ Malczewski41[Uncaptioned image], S. Malde64[Uncaptioned image], L. Malentacca49[Uncaptioned image], A. Malinin44[Uncaptioned image], T. Maltsev44[Uncaptioned image], G. Manca32,k[Uncaptioned image], G. Mancinelli13[Uncaptioned image], C. Mancuso30,14,n[Uncaptioned image], R. Manera Escalero45[Uncaptioned image], F. M.  Manganella37[Uncaptioned image], D. Manuzzi25[Uncaptioned image], D. Marangotto30,n[Uncaptioned image], J.F. Marchand10[Uncaptioned image], R. Marchevski50[Uncaptioned image], U. Marconi25[Uncaptioned image], E. Mariani16[Uncaptioned image], S. Mariani49[Uncaptioned image], C. Marin Benito45,49[Uncaptioned image], J. Marks22[Uncaptioned image], A.M. Marshall55[Uncaptioned image], L.  Martel64[Uncaptioned image], G. Martelli34,r[Uncaptioned image], G. Martellotti36[Uncaptioned image], L. Martinazzoli49[Uncaptioned image], M. Martinelli31,o[Uncaptioned image], D.  Martinez Gomez78[Uncaptioned image], D. Martinez Santos81[Uncaptioned image], F. Martinez Vidal48[Uncaptioned image], A.  Martorell i Granollers46[Uncaptioned image], A. Massafferri2[Uncaptioned image], R. Matev49[Uncaptioned image], A. Mathad49[Uncaptioned image], V. Matiunin44[Uncaptioned image], C. Matteuzzi69[Uncaptioned image], K.R. Mattioli15[Uncaptioned image], A. Mauri62[Uncaptioned image], E. Maurice15[Uncaptioned image], J. Mauricio45[Uncaptioned image], P. Mayencourt50[Uncaptioned image], J. Mazorra de Cos48[Uncaptioned image], M. Mazurek42[Uncaptioned image], M. McCann62[Uncaptioned image], L. Mcconnell23[Uncaptioned image], T.H. McGrath63[Uncaptioned image], N.T. McHugh60[Uncaptioned image], A. McNab63[Uncaptioned image], R. McNulty14,23[Uncaptioned image], B. Meadows66[Uncaptioned image], G. Meier19[Uncaptioned image], D. Melnychuk42[Uncaptioned image], F. M.  Meng4,b[Uncaptioned image], M. Merk38,79[Uncaptioned image], A. Merli50[Uncaptioned image], L. Meyer Garcia67[Uncaptioned image], D. Miao5,7[Uncaptioned image], H. Miao7[Uncaptioned image], M. Mikhasenko76[Uncaptioned image], D.A. Milanes75[Uncaptioned image], A. Minotti31,o[Uncaptioned image], E. Minucci28[Uncaptioned image], T. Miralles11[Uncaptioned image], B. Mitreska19[Uncaptioned image], D.S. Mitzel19[Uncaptioned image], A. Modak58[Uncaptioned image], R.A. Mohammed64[Uncaptioned image], R.D. Moise17[Uncaptioned image], S. Mokhnenko44[Uncaptioned image], E. F. Molina Cardenas83[Uncaptioned image], T. Mombächer49[Uncaptioned image], M. Monk57,1[Uncaptioned image], S. Monteil11[Uncaptioned image], A. Morcillo Gomez47[Uncaptioned image], G. Morello28[Uncaptioned image], M.J. Morello35,s[Uncaptioned image], M.P. Morgenthaler22[Uncaptioned image], J. Moron40[Uncaptioned image], W.  Morren38[Uncaptioned image], A.B. Morris49[Uncaptioned image], A.G. Morris13[Uncaptioned image], R. Mountain69[Uncaptioned image], H. Mu4,b[Uncaptioned image], Z. M.  Mu6[Uncaptioned image], E. Muhammad57[Uncaptioned image], F. Muheim59[Uncaptioned image], M. Mulder78[Uncaptioned image], K. Müller51[Uncaptioned image], F. Muñoz-Rojas9[Uncaptioned image], R. Murta62[Uncaptioned image], P. Naik61[Uncaptioned image], T. Nakada50[Uncaptioned image], R. Nandakumar58[Uncaptioned image], T. Nanut49[Uncaptioned image], I. Nasteva3[Uncaptioned image], M. Needham59[Uncaptioned image], N. Neri30,n[Uncaptioned image], S. Neubert18[Uncaptioned image], N. Neufeld49[Uncaptioned image], P. Neustroev44, J. Nicolini19,14[Uncaptioned image], D. Nicotra79[Uncaptioned image], E.M. Niel49[Uncaptioned image], N. Nikitin44[Uncaptioned image], Q. Niu73[Uncaptioned image], P. Nogarolli3[Uncaptioned image], P. Nogga18[Uncaptioned image], C. Normand55[Uncaptioned image], J. Novoa Fernandez47[Uncaptioned image], G. Nowak66[Uncaptioned image], C. Nunez83[Uncaptioned image], H. N.  Nur60[Uncaptioned image], A. Oblakowska-Mucha40[Uncaptioned image], V. Obraztsov44[Uncaptioned image], T. Oeser17[Uncaptioned image], S. Okamura26,l[Uncaptioned image], A. Okhotnikov44, O. Okhrimenko53[Uncaptioned image], R. Oldeman32,k[Uncaptioned image], F. Oliva59[Uncaptioned image], M. Olocco19[Uncaptioned image], C.J.G. Onderwater79[Uncaptioned image], R.H. O’Neil59[Uncaptioned image], D. Osthues19[Uncaptioned image], J.M. Otalora Goicochea3[Uncaptioned image], P. Owen51[Uncaptioned image], A. Oyanguren48[Uncaptioned image], O. Ozcelik59[Uncaptioned image], F. Paciolla35,w[Uncaptioned image], A.  Padee42[Uncaptioned image], K.O. Padeken18[Uncaptioned image], B. Pagare57[Uncaptioned image], P.R. Pais22[Uncaptioned image], T. Pajero49[Uncaptioned image], A. Palano24[Uncaptioned image], M. Palutan28[Uncaptioned image], X.  Pan4,b[Uncaptioned image], G. Panshin44[Uncaptioned image], L. Paolucci57[Uncaptioned image], A. Papanestis58,49[Uncaptioned image], M. Pappagallo24,h[Uncaptioned image], $L.$ Pappalardo26,l[Uncaptioned image], C. Pappenheimer66[Uncaptioned image], C. Parkes63[Uncaptioned image], D.  Parmar76[Uncaptioned image], B. Passalacqua26,l[Uncaptioned image], G. Passaleva27[Uncaptioned image], D. Passaro35,s[Uncaptioned image], A. Pastore24[Uncaptioned image], M. Patel62[Uncaptioned image], J. Patoc64[Uncaptioned image], C. Patrignani25,j[Uncaptioned image], A.  Paul69[Uncaptioned image], C.J. Pawley79[Uncaptioned image], A. Pellegrino38[Uncaptioned image], J.  Peng5,7[Uncaptioned image], M. Pepe Altarelli28[Uncaptioned image], S. Perazzini25[Uncaptioned image], D. Pereima44[Uncaptioned image], H.  Pereira Da Costa68[Uncaptioned image], A. Pereiro Castro47[Uncaptioned image], P. Perret11[Uncaptioned image], A.  Perrevoort78[Uncaptioned image], A. Perro49,13[Uncaptioned image], M.J. Peters66[Uncaptioned image], K. Petridis55[Uncaptioned image], A. Petrolini29,m[Uncaptioned image], J. P.  Pfaller66[Uncaptioned image], H. Pham69[Uncaptioned image], L. Pica35,s[Uncaptioned image], M. Piccini34[Uncaptioned image], L.  Piccolo32[Uncaptioned image], B. Pietrzyk10[Uncaptioned image], G. Pietrzyk14[Uncaptioned image], D. Pinci36[Uncaptioned image], F. Pisani49[Uncaptioned image], M. Pizzichemi31,o,49[Uncaptioned image], V. M. Placinta43[Uncaptioned image], M. Plo Casasus47[Uncaptioned image], T. Poeschl49[Uncaptioned image], F. Polci16,49[Uncaptioned image], M. Poli Lener28[Uncaptioned image], A. Poluektov13[Uncaptioned image], N. Polukhina44[Uncaptioned image], I. Polyakov44[Uncaptioned image], E. Polycarpo3[Uncaptioned image], S. Ponce49[Uncaptioned image], D. Popov7[Uncaptioned image], S. Poslavskii44[Uncaptioned image], K. Prasanth59[Uncaptioned image], C. Prouve81[Uncaptioned image], D. Provenzano32,k[Uncaptioned image], V. Pugatch53[Uncaptioned image], G. Punzi35,t[Uncaptioned image], S.  Qasim51[Uncaptioned image], Q. Q.  Qian6[Uncaptioned image], W. Qian7[Uncaptioned image], N. Qin4,b[Uncaptioned image], S. Qu4,b[Uncaptioned image], R. Quagliani49[Uncaptioned image], R.I. Rabadan Trejo57[Uncaptioned image], J.H. Rademacker55[Uncaptioned image], M. Rama35[Uncaptioned image], M.  Ramírez García83[Uncaptioned image], V. Ramos De Oliveira70[Uncaptioned image], M. Ramos Pernas57[Uncaptioned image], M.S. Rangel3[Uncaptioned image], F. Ratnikov44[Uncaptioned image], G. Raven39[Uncaptioned image], M. Rebollo De Miguel48[Uncaptioned image], F. Redi30,i[Uncaptioned image], J. Reich55[Uncaptioned image], F. Reiss63[Uncaptioned image], Z. Ren7[Uncaptioned image], P.K. Resmi64[Uncaptioned image], R. Ribatti50[Uncaptioned image], G. Ricart15,12[Uncaptioned image], D. Riccardi35,s[Uncaptioned image], S. Ricciardi58[Uncaptioned image], K. Richardson65[Uncaptioned image], M. Richardson-Slipper59[Uncaptioned image], K. Rinnert61[Uncaptioned image], P. Robbe14,49[Uncaptioned image], G. Robertson60[Uncaptioned image], E. Rodrigues61[Uncaptioned image], A. Rodriguez Alvarez45[Uncaptioned image], E. Rodriguez Fernandez47[Uncaptioned image], J.A. Rodriguez Lopez75[Uncaptioned image], E. Rodriguez Rodriguez47[Uncaptioned image], J. Roensch19[Uncaptioned image], A. Rogachev44[Uncaptioned image], A. Rogovskiy58[Uncaptioned image], D.L. Rolf49[Uncaptioned image], P. Roloff49[Uncaptioned image], V. Romanovskiy66[Uncaptioned image], A. Romero Vidal47[Uncaptioned image], G. Romolini26[Uncaptioned image], F. Ronchetti50[Uncaptioned image], T. Rong6[Uncaptioned image], M. Rotondo28[Uncaptioned image], S. R.  Roy22[Uncaptioned image], M.S. Rudolph69[Uncaptioned image], M. Ruiz Diaz22[Uncaptioned image], R.A. Ruiz Fernandez47[Uncaptioned image], J. Ruiz Vidal82,aa[Uncaptioned image], A. Ryzhikov44[Uncaptioned image], J. Ryzka40[Uncaptioned image], J. J. Saavedra-Arias9[Uncaptioned image], J.J. Saborido Silva47[Uncaptioned image], R. Sadek15[Uncaptioned image], N. Sagidova44[Uncaptioned image], D. Sahoo77[Uncaptioned image], N. Sahoo54[Uncaptioned image], B. Saitta32,k[Uncaptioned image], M. Salomoni31,49,o[Uncaptioned image], I. Sanderswood48[Uncaptioned image], R. Santacesaria36[Uncaptioned image], C. Santamarina Rios47[Uncaptioned image], M. Santimaria28,49[Uncaptioned image], L. Santoro 2[Uncaptioned image], E. Santovetti37[Uncaptioned image], A. Saputi26,49[Uncaptioned image], D. Saranin44[Uncaptioned image], A. Sarnatskiy78[Uncaptioned image], G. Sarpis59[Uncaptioned image], M. Sarpis63[Uncaptioned image], C. Satriano36,u[Uncaptioned image], A. Satta37[Uncaptioned image], M. Saur6[Uncaptioned image], D. Savrina44[Uncaptioned image], H. Sazak17[Uncaptioned image], F. Sborzacchi49,28[Uncaptioned image], L.G. Scantlebury Smead64[Uncaptioned image], A. Scarabotto19[Uncaptioned image], S. Schael17[Uncaptioned image], S. Scherl61[Uncaptioned image], M. Schiller60[Uncaptioned image], H. Schindler49[Uncaptioned image], M. Schmelling21[Uncaptioned image], B. Schmidt49[Uncaptioned image], S. Schmitt17[Uncaptioned image], H. Schmitz18, O. Schneider50[Uncaptioned image], A. Schopper49[Uncaptioned image], N. Schulte19[Uncaptioned image], S. Schulte50[Uncaptioned image], M.H. Schune14[Uncaptioned image], R. Schwemmer49[Uncaptioned image], G. Schwering17[Uncaptioned image], B. Sciascia28[Uncaptioned image], A. Sciuccati49[Uncaptioned image], I. Segal76[Uncaptioned image], S. Sellam47[Uncaptioned image], A. Semennikov44[Uncaptioned image], T. Senger51[Uncaptioned image], M. Senghi Soares39[Uncaptioned image], A. Sergi29,m[Uncaptioned image], N. Serra51[Uncaptioned image], L. Sestini33[Uncaptioned image], A. Seuthe19[Uncaptioned image], Y. Shang6[Uncaptioned image], D.M. Shangase83[Uncaptioned image], M. Shapkin44[Uncaptioned image], R. S.  Sharma69[Uncaptioned image], I. Shchemerov44[Uncaptioned image], L. Shchutska50[Uncaptioned image], T. Shears61[Uncaptioned image], L. Shekhtman44[Uncaptioned image], Z. Shen6[Uncaptioned image], S. Sheng5,7[Uncaptioned image], V. Shevchenko44[Uncaptioned image], B. Shi7[Uncaptioned image], Q. Shi7[Uncaptioned image], Y. Shimizu14[Uncaptioned image], E. Shmanin25[Uncaptioned image], R. Shorkin44[Uncaptioned image], J.D. Shupperd69[Uncaptioned image], R. Silva Coutinho69[Uncaptioned image], G. Simi33,q[Uncaptioned image], S. Simone24,h[Uncaptioned image], N. Skidmore57[Uncaptioned image], T. Skwarnicki69[Uncaptioned image], M.W. Slater54[Uncaptioned image], J.C. Smallwood64[Uncaptioned image], E. Smith65[Uncaptioned image], K. Smith68[Uncaptioned image], M. Smith62[Uncaptioned image], A. Snoch38[Uncaptioned image], L. Soares Lavra59[Uncaptioned image], M.D. Sokoloff66[Uncaptioned image], F.J.P. Soler60[Uncaptioned image], A. Solomin44,55[Uncaptioned image], A. Solovev44[Uncaptioned image], I. Solovyev44[Uncaptioned image], N. S.  Sommerfeld18[Uncaptioned image], R. Song1[Uncaptioned image], Y. Song50[Uncaptioned image], Y. Song4,b[Uncaptioned image], Y. S.  Song6[Uncaptioned image], F.L. Souza De Almeida69[Uncaptioned image], B. Souza De Paula3[Uncaptioned image], E. Spadaro Norella29,m[Uncaptioned image], E. Spedicato25[Uncaptioned image], J.G. Speer19[Uncaptioned image], E. Spiridenkov44, P. Spradlin60[Uncaptioned image], V. Sriskaran49[Uncaptioned image], F. Stagni49[Uncaptioned image], M. Stahl49[Uncaptioned image], S. Stahl49[Uncaptioned image], S. Stanislaus64[Uncaptioned image], E.N. Stein49[Uncaptioned image], O. Steinkamp51[Uncaptioned image], O. Stenyakin44, H. Stevens19[Uncaptioned image], D. Strekalina44[Uncaptioned image], Y. Su7[Uncaptioned image], F. Suljik64[Uncaptioned image], J. Sun32[Uncaptioned image], L. Sun74[Uncaptioned image], D. Sundfeld2[Uncaptioned image], W. Sutcliffe51[Uncaptioned image], P.N. Swallow54[Uncaptioned image], K. Swientek40[Uncaptioned image], F. Swystun56[Uncaptioned image], A. Szabelski42[Uncaptioned image], T. Szumlak40[Uncaptioned image], Y. Tan4,b[Uncaptioned image], Y. Tang74[Uncaptioned image], M.D. Tat64[Uncaptioned image], A. Terentev44[Uncaptioned image], F. Terzuoli35,w,49[Uncaptioned image], F. Teubert49[Uncaptioned image], E. Thomas49[Uncaptioned image], D.J.D. Thompson54[Uncaptioned image], H. Tilquin62[Uncaptioned image], V. Tisserand11[Uncaptioned image], S. T’Jampens10[Uncaptioned image], M. Tobin5,49[Uncaptioned image], L. Tomassetti26,l[Uncaptioned image], G. Tonani30,n,49[Uncaptioned image], X. Tong6[Uncaptioned image], D. Torres Machado2[Uncaptioned image], L. Toscano19[Uncaptioned image], D.Y. Tou4,b[Uncaptioned image], C. Trippl46[Uncaptioned image], G. Tuci22[Uncaptioned image], N. Tuning38[Uncaptioned image], L.H. Uecker22[Uncaptioned image], A. Ukleja40[Uncaptioned image], D.J. Unverzagt22[Uncaptioned image], B.  Urbach59[Uncaptioned image], E. Ursov44[Uncaptioned image], A. Usachov39[Uncaptioned image], A. Ustyuzhanin44[Uncaptioned image], U. Uwer22[Uncaptioned image], V. Vagnoni25[Uncaptioned image], V.  Valcarce Cadenas47[Uncaptioned image], G. Valenti25[Uncaptioned image], N. Valls Canudas49[Uncaptioned image], H. Van Hecke68[Uncaptioned image], E. van Herwijnen62[Uncaptioned image], C.B. Van Hulse47,y[Uncaptioned image], R. Van Laak50[Uncaptioned image], M. van Veghel38[Uncaptioned image], G. Vasquez51[Uncaptioned image], R. Vazquez Gomez45[Uncaptioned image], P. Vazquez Regueiro47[Uncaptioned image], C. Vázquez Sierra47[Uncaptioned image], S. Vecchi26[Uncaptioned image], $J.$ Velthuis55[Uncaptioned image], M. Veltri27,x[Uncaptioned image], A. Venkateswaran50[Uncaptioned image], M. Verdoglia32[Uncaptioned image], M. Vesterinen57[Uncaptioned image], D.  Vico Benet64[Uncaptioned image], P.  Vidrier Villalba45[Uncaptioned image], M. Vieites Diaz49[Uncaptioned image], X. Vilasis-Cardona46[Uncaptioned image], E. Vilella Figueras61[Uncaptioned image], A. Villa25[Uncaptioned image], P. Vincent16[Uncaptioned image], F.C. Volle54[Uncaptioned image], D. vom Bruch13[Uncaptioned image], N. Voropaev44[Uncaptioned image], K. Vos79[Uncaptioned image], C. Vrahas59[Uncaptioned image], J. Wagner19[Uncaptioned image], J. Walsh35[Uncaptioned image], E.J. Walton1,57[Uncaptioned image], G. Wan6[Uncaptioned image], C. Wang22[Uncaptioned image], G. Wang8[Uncaptioned image], H. Wang73[Uncaptioned image], J. Wang6[Uncaptioned image], J. Wang5[Uncaptioned image], J. Wang4,b[Uncaptioned image], J. Wang74[Uncaptioned image], M. Wang30[Uncaptioned image], N. W.  Wang7[Uncaptioned image], R. Wang55[Uncaptioned image], X. Wang8[Uncaptioned image], X. Wang72[Uncaptioned image], X. W.  Wang62[Uncaptioned image], Y. Wang6[Uncaptioned image], Y. W.  Wang73[Uncaptioned image], Z. Wang14[Uncaptioned image], Z. Wang4,b[Uncaptioned image], Z. Wang30[Uncaptioned image], J.A. Ward57,1[Uncaptioned image], M. Waterlaat49[Uncaptioned image], N.K. Watson54[Uncaptioned image], D. Websdale62[Uncaptioned image], Y. Wei6[Uncaptioned image], J. Wendel81[Uncaptioned image], B.D.C. Westhenry55[Uncaptioned image], C. White56[Uncaptioned image], M. Whitehead60[Uncaptioned image], E. Whiter54[Uncaptioned image], A.R. Wiederhold63[Uncaptioned image], D. Wiedner19[Uncaptioned image], G. Wilkinson64[Uncaptioned image], M.K. Wilkinson66[Uncaptioned image], M. Williams65[Uncaptioned image], M. J. Williams49[Uncaptioned image], M.R.J. Williams59[Uncaptioned image], R. Williams56[Uncaptioned image], Z.  Williams55[Uncaptioned image], $F.$ Wilson58[Uncaptioned image], M. Winn12[Uncaptioned image], W. Wislicki42[Uncaptioned image], M. Witek41[Uncaptioned image], L. Witola22[Uncaptioned image], G. Wormser14[Uncaptioned image], S.A. Wotton56[Uncaptioned image], H. Wu69[Uncaptioned image], J. Wu8[Uncaptioned image], X. Wu74[Uncaptioned image], Y. Wu6[Uncaptioned image], Z. Wu7[Uncaptioned image], K. Wyllie49[Uncaptioned image], S. Xian72[Uncaptioned image], Z. Xiang5[Uncaptioned image], Y. Xie8[Uncaptioned image], A. Xu35[Uncaptioned image], J. Xu7[Uncaptioned image], L. Xu4,b[Uncaptioned image], L. Xu4,b[Uncaptioned image], M. Xu57[Uncaptioned image], Z. Xu49[Uncaptioned image], Z. Xu7[Uncaptioned image], Z. Xu5[Uncaptioned image], K.  Yang62[Uncaptioned image], S. Yang7[Uncaptioned image], X. Yang6[Uncaptioned image], Y. Yang29,m[Uncaptioned image], Z. Yang6[Uncaptioned image], V. Yeroshenko14[Uncaptioned image], H. Yeung63[Uncaptioned image], H. Yin8[Uncaptioned image], X.  Yin7[Uncaptioned image], C. Y.  Yu6[Uncaptioned image], J. Yu71[Uncaptioned image], X. Yuan5[Uncaptioned image], Y Yuan5,7[Uncaptioned image], E. Zaffaroni50[Uncaptioned image], M. Zavertyaev21[Uncaptioned image], M. Zdybal41[Uncaptioned image], F. Zenesini25,j[Uncaptioned image], C.  Zeng5,7[Uncaptioned image], M. Zeng4,b[Uncaptioned image], C. Zhang6[Uncaptioned image], D. Zhang8[Uncaptioned image], J. Zhang7[Uncaptioned image], L. Zhang4,b[Uncaptioned image], S. Zhang71[Uncaptioned image], S. Zhang64[Uncaptioned image], Y. Zhang6[Uncaptioned image], Y. Z.  Zhang4,b[Uncaptioned image], Z. Zhang4,b[Uncaptioned image], Y. Zhao22[Uncaptioned image], A. Zharkova44[Uncaptioned image], A. Zhelezov22[Uncaptioned image], S. Z.  Zheng6[Uncaptioned image], X. Z.  Zheng4,b[Uncaptioned image], Y. Zheng7[Uncaptioned image], T. Zhou6[Uncaptioned image], X. Zhou8[Uncaptioned image], Y. Zhou7[Uncaptioned image], V. Zhovkovska57[Uncaptioned image], L. Z.  Zhu7[Uncaptioned image], X. Zhu4,b[Uncaptioned image], X. Zhu8[Uncaptioned image], V. Zhukov17[Uncaptioned image], J. Zhuo48[Uncaptioned image], Q. Zou5,7[Uncaptioned image], D. Zuliani33,q[Uncaptioned image], G. Zunica50[Uncaptioned image].
1School of Physics and Astronomy, Monash University, Melbourne, Australia
2Centro Brasileiro de Pesquisas Físicas (CBPF), Rio de Janeiro, Brazil
3Universidade Federal do Rio de Janeiro (UFRJ), Rio de Janeiro, Brazil
4Department of Engineering Physics, Tsinghua University, Beijing, China
5Institute Of High Energy Physics (IHEP), Beijing, China
6School of Physics State Key Laboratory of Nuclear Physics and Technology, Peking University, Beijing, China
7University of Chinese Academy of Sciences, Beijing, China
8Institute of Particle Physics, Central China Normal University, Wuhan, Hubei, China
9Consejo Nacional de Rectores (CONARE), San Jose, Costa Rica
10Université Savoie Mont Blanc, CNRS, IN2P3-LAPP, Annecy, France
11Université Clermont Auvergne, CNRS/IN2P3, LPC, Clermont-Ferrand, France
12Université Paris-Saclay, Centre d’Etudes de Saclay (CEA), IRFU, Saclay, France, Gif-Sur-Yvette, France
13Aix Marseille Univ, CNRS/IN2P3, CPPM, Marseille, France
14Université Paris-Saclay, CNRS/IN2P3, IJCLab, Orsay, France
15Laboratoire Leprince-Ringuet, CNRS/IN2P3, Ecole Polytechnique, Institut Polytechnique de Paris, Palaiseau, France
16LPNHE, Sorbonne Université, Paris Diderot Sorbonne Paris Cité, CNRS/IN2P3, Paris, France
17I. Physikalisches Institut, RWTH Aachen University, Aachen, Germany
18Universität Bonn - Helmholtz-Institut für Strahlen und Kernphysik, Bonn, Germany
19Fakultät Physik, Technische Universität Dortmund, Dortmund, Germany
20Physikalisches Institut, Albert-Ludwigs-Universität Freiburg, Freiburg, Germany
21Max-Planck-Institut für Kernphysik (MPIK), Heidelberg, Germany
22Physikalisches Institut, Ruprecht-Karls-Universität Heidelberg, Heidelberg, Germany
23School of Physics, University College Dublin, Dublin, Ireland
24INFN Sezione di Bari, Bari, Italy
25INFN Sezione di Bologna, Bologna, Italy
26INFN Sezione di Ferrara, Ferrara, Italy
27INFN Sezione di Firenze, Firenze, Italy
28INFN Laboratori Nazionali di Frascati, Frascati, Italy
29INFN Sezione di Genova, Genova, Italy
30INFN Sezione di Milano, Milano, Italy
31INFN Sezione di Milano-Bicocca, Milano, Italy
32INFN Sezione di Cagliari, Monserrato, Italy
33INFN Sezione di Padova, Padova, Italy
34INFN Sezione di Perugia, Perugia, Italy
35INFN Sezione di Pisa, Pisa, Italy
36INFN Sezione di Roma La Sapienza, Roma, Italy
37INFN Sezione di Roma Tor Vergata, Roma, Italy
38Nikhef National Institute for Subatomic Physics, Amsterdam, Netherlands
39Nikhef National Institute for Subatomic Physics and VU University Amsterdam, Amsterdam, Netherlands
40AGH - University of Krakow, Faculty of Physics and Applied Computer Science, Kraków, Poland
41Henryk Niewodniczanski Institute of Nuclear Physics Polish Academy of Sciences, Kraków, Poland
42National Center for Nuclear Research (NCBJ), Warsaw, Poland
43Horia Hulubei National Institute of Physics and Nuclear Engineering, Bucharest-Magurele, Romania
44Authors affiliated with an institute formerly covered by a cooperation agreement with CERN.
45ICCUB, Universitat de Barcelona, Barcelona, Spain
46La Salle, Universitat Ramon Llull, Barcelona, Spain
47Instituto Galego de Física de Altas Enerxías (IGFAE), Universidade de Santiago de Compostela, Santiago de Compostela, Spain
48Instituto de Fisica Corpuscular, Centro Mixto Universidad de Valencia - CSIC, Valencia, Spain
49European Organization for Nuclear Research (CERN), Geneva, Switzerland
50Institute of Physics, Ecole Polytechnique Fédérale de Lausanne (EPFL), Lausanne, Switzerland
51Physik-Institut, Universität Zürich, Zürich, Switzerland
52NSC Kharkiv Institute of Physics and Technology (NSC KIPT), Kharkiv, Ukraine
53Institute for Nuclear Research of the National Academy of Sciences (KINR), Kyiv, Ukraine
54School of Physics and Astronomy, University of Birmingham, Birmingham, United Kingdom
55H.H. Wills Physics Laboratory, University of Bristol, Bristol, United Kingdom
56Cavendish Laboratory, University of Cambridge, Cambridge, United Kingdom
57Department of Physics, University of Warwick, Coventry, United Kingdom
58STFC Rutherford Appleton Laboratory, Didcot, United Kingdom
59School of Physics and Astronomy, University of Edinburgh, Edinburgh, United Kingdom
60School of Physics and Astronomy, University of Glasgow, Glasgow, United Kingdom
61Oliver Lodge Laboratory, University of Liverpool, Liverpool, United Kingdom
62Imperial College London, London, United Kingdom
63Department of Physics and Astronomy, University of Manchester, Manchester, United Kingdom
64Department of Physics, University of Oxford, Oxford, United Kingdom
65Massachusetts Institute of Technology, Cambridge, MA, United States
66University of Cincinnati, Cincinnati, OH, United States
67University of Maryland, College Park, MD, United States
68Los Alamos National Laboratory (LANL), Los Alamos, NM, United States
69Syracuse University, Syracuse, NY, United States
70Pontifícia Universidade Católica do Rio de Janeiro (PUC-Rio), Rio de Janeiro, Brazil, associated to 3
71School of Physics and Electronics, Hunan University, Changsha City, China, associated to 8
72Guangdong Provincial Key Laboratory of Nuclear Science, Guangdong-Hong Kong Joint Laboratory of Quantum Matter, Institute of Quantum Matter, South China Normal University, Guangzhou, China, associated to 4
73Lanzhou University, Lanzhou, China, associated to 5
74School of Physics and Technology, Wuhan University, Wuhan, China, associated to 4
75Departamento de Fisica , Universidad Nacional de Colombia, Bogota, Colombia, associated to 16
76Ruhr Universitaet Bochum, Fakultaet f. Physik und Astronomie, Bochum, Germany, associated to 19
77Eotvos Lorand University, Budapest, Hungary, associated to 49
78Van Swinderen Institute, University of Groningen, Groningen, Netherlands, associated to 38
79Universiteit Maastricht, Maastricht, Netherlands, associated to 38
80Tadeusz Kosciuszko Cracow University of Technology, Cracow, Poland, associated to 41
81Universidade da Coruña, A Coruña, Spain, associated to 46
82Department of Physics and Astronomy, Uppsala University, Uppsala, Sweden, associated to 60
83University of Michigan, Ann Arbor, MI, United States, associated to 69
aCentro Federal de Educacão Tecnológica Celso Suckow da Fonseca, Rio De Janeiro, Brazil
bCenter for High Energy Physics, Tsinghua University, Beijing, China
cHangzhou Institute for Advanced Study, UCAS, Hangzhou, China
dSchool of Physics and Electronics, Henan University , Kaifeng, China
eLIP6, Sorbonne Université, Paris, France
fLamarr Institute for Machine Learning and Artificial Intelligence, Dortmund, Germany
gUniversidad Nacional Autónoma de Honduras, Tegucigalpa, Honduras
hUniversità di Bari, Bari, Italy
iUniversità di Bergamo, Bergamo, Italy
jUniversità di Bologna, Bologna, Italy
kUniversità di Cagliari, Cagliari, Italy
lUniversità di Ferrara, Ferrara, Italy
mUniversità di Genova, Genova, Italy
nUniversità degli Studi di Milano, Milano, Italy
oUniversità degli Studi di Milano-Bicocca, Milano, Italy
pUniversità di Modena e Reggio Emilia, Modena, Italy
qUniversità di Padova, Padova, Italy
rUniversità di Perugia, Perugia, Italy
sScuola Normale Superiore, Pisa, Italy
tUniversità di Pisa, Pisa, Italy
uUniversità della Basilicata, Potenza, Italy
vUniversità di Roma Tor Vergata, Roma, Italy
wUniversità di Siena, Siena, Italy
xUniversità di Urbino, Urbino, Italy
yUniversidad de Alcalá, Alcalá de Henares , Spain
zFacultad de Ciencias Fisicas, Madrid, Spain
aaDepartment of Physics/Division of Particle Physics, Lund, Sweden
Deceased